### v1

# Gravity versus astrophysics in black hole images and photon rings: equatorial emissions and spherically symmetric spacetimes

Abstract
Context. The Event Horizon Telescope collaboration released in 2019 the first horizon-scale images of a black hole accretion flow, opening a novel route for plasma physics comprehension and gravitational tests. Although the present unresolved images deeply depend on the astrophysical properties of the accreted matter, general relativity predicts that they contain highly lensed observables, the so-called ‘photon rings’, embodying the effects of strong-field gravity.
Aims. Focusing on the particular case of the supermassive black hole M87* and adopting a geometrically thin, equatorial disc as a phenomenological configuration for the accreting matter, our goal is to study the degeneracy of spacetime curvature and of physically-motivated emission processes on plane-of-sky EHT-like images observed at 230 and 345 GHz.
Methods. In a parametric framework, we simulate adaptively ray-traced images using the code GYOTO in various spherically-symmetric spacetime geometries, for a comprehensive class of disc velocities and a library of realistic synchrotron emission profiles. We then extract the width and the peak position of 1D intensity cross sections on the direct image and the first photon ring.
Results. We show that, among the investigated quantities, the most appropriate observables to probe the geometry are the peak positions of the first photon ring. Small geometric deviations can be unequivocally detected, regardless of the motion of the disc, ranging from a Keplerian to a radially infalling one, if the black hole mass-to-distance estimate is accurate up to around 2%; the current uncertainty of 11% being sufficient just to access extreme deviations.
Conclusions. The equatorial set-up of this paper, favoured by present EHT observations of M87*, is adapted to model future measurements at higher observing frequencies, where absorption effects are negligible, and with higher resolution, indispensable to resolve the photon rings. Additional work is needed to investigate if our conclusions hold for more realistic disc configurations.
Key Words.:
gravitation – accretion, accretion discs – black hole physics – relativistic processes – galaxies: individual: M87
1 Introduction
The existence of black holes (BHs) is a direct consequence of general relativity (GR), the standard theory of gravity, and their first solution was found by Karl Schwarzshild (Schwarzschild 1916) a few weeks after the publication of the GR foundations. A cardinal generalisation to rotating black holes, published by Roy Kerr (Kerr 1963), represents the only [^1] With the astrophysically pertinent assumption of null electric charge. stationary solution satisfying the hypotheses of the no-hair theorem (Israel 1967; Carter 1971; Hawking 1972; Robinson 1975) and a Kerr black hole is entirely described by its mass and angular momentum.
The definition of a black hole relies on the presence of an immaterial boundary, the event horizon, that can be crossed only in one direction, towards the BH: not even light can escape it. Also, according to the cosmic censorship conjecture (Penrose 1969), for a Kerr or a Schwarzschild black hole, the event horizon hides observationally inaccessible incomplete null geodesics that may be associated with a curvature singularity. On top of that, the interior of the Kerr spacetime contains closed timelike curves that correspond to a causality violation (Carter 1968). These predictions are controversial and they might mark the limit of the applicability domain of general relativity; it is then justified to continue testing this centennial theory, already thoroughly corroborated for weak gravitational fields, in the extreme conditions of the strong-field regime begotten by black holes.
Since a black hole is not a luminous source itself, its presence is inferred from the influence it has on its environment, which has been revealed by recent revolutionary astronomical observations. The first detection of a gravitational wave signal from the merger of two black holes was announced in 2016 by the LIGO-Virgo collaboration (LIGO Scientific Collaboration and Virgo Collaboration 2016). Two other great achievements, both by GRAVITY in 2018, were the astrometric measurements, close to the supermassive black hole (SMBH) Sagittarius A*, of the orbital motion of the S stars (GRAVITY collaboration 2018a) and of the high states, or flares, of variable near-infrared emission (GRAVITY collaboration 2018b). Finally, the first breakthrough images of black hole accretion flows, probing the strong-field gravity region just above the event horizon, has been released by the Event Horizon Telescope (EHT) in 2019 for M87* (The Event Horizon Telescope Collaboration 2019a).
These images, although compatible with the GR-predicted observational appearance of a Kerr black hole, do not exclude alternative spacetimes (The Event Horizon Telescope Collaboration 2019d; Vincent et al. 2021; Gralla 2021). However, consistency tests of the Kerr-hypothesis will become more stringent with the upgrades of the Next Generation EHT (ngEHT, Johnson et al. 2023; Doeleman et al. 2023; Ayzenberg et al. 2023) and, particularly, with future space-based very-long-baseline interferometers like the proposed mission Black Hole Explorer (BHEX, Johnson et al. 2024; Hudson et al. 2023), that will achieve sufficient angular resolution to access the most gravitationally affected features of the image, namely the photon rings.
It has already been widely shown that the visual aspect of the accretion flow depends both on the spacetime geometry and the properties of the emission, and that the observational appearance of M87*, under the limited instrumental resolution of the EHT, can be mimicked by non-Kerr objects (for instance, Vincent et al. 2021). In this paper we do not attempt to model the limitations of any particular existing or future observing instrument, assuming that a perfect image-domain reconstruction can be obtained. Thus, we focus on the degeneracy between gravity and astrophysics in the images that is fundamental and theoretical in nature. Hence, it is crucial to identify observable components of the images which allow to disentangle these effects. In this work we focus on photon rings, that is, lensed secondary images of the accretion disc superposed on top of the primary image and formed by photons executing at least half-loop around the black hole. In contrast to the primary image and the central depression in brightness, known as the ‘observable shadow’, which are strongly dependent on the astrophysical configuration (The Event Horizon Telescope Collaboration 2019c; Bauer et al. 2022; Kocherlakota & Rezzolla 2022), photon rings are predominantly impacted by the spacetime geometry (Desire et al. 2025).
Photon rings are labelled by the number $n$ of half turns of the photons around the central body. It has been shown that they follow a specific functional form for the Kerr metric not only at high-$n$ (Gralla et al. 2020), but also at $n=1$ (Cárdenas-Avendaño & Lupsasca 2023). Besides, as shown in the latter paper for $n=1$ and elsewhere for $n\geq 2$ (Paugnat et al. 2022), the corresponding interferometric signature is robust under change of the astrophysical configuration for a Kerr black hole. A limitation present in these works, that we would like to overcome here, is the adoption of a three-parameter subset of Johnson’s Standard-Unbounded distribution (Gralla et al. 2020) without physical parameters describing the emission process.
Properties of the photon rings have been used not only to carry out the consistency tests mentioned in the previous paragraph, but also to investigate discriminatory tests of GR, at least for specific classes of metrics and heuristic emission processes:
  • Bauer et al. (2022) and Kocherlakota & Rezzolla (2022) both investigated the properties of the intensity collected on EHT-like images of spherically-symmetric static compact objects surrounded by spherically-infalling emitting matter. Thin-disc models were also studied, as in Eichhorn et al. (2023);
  • Several authors suggested to use Lyapunov exponents to differentiate between spacetime metrics of compact objects that admit separable geodesic equations (Wielgus 2021; Staelens et al. 2023; da Silva et al. 2023; Kocherlakota et al. 2024b). However, since that theoretical feature of photon rings is not directly detectable, one should invoke, for instance, their link with potentially measurable brightness autocorrelations (Hadar et al. 2021; Chesler et al. 2021).
Finally, a promising perspective for the study of photon rings consists in the analysis of their polarised signal (Himwich et al. 2020), which could lead to the detection of the first photon ring before reaching the baselines needed to resolve its interferometric imprint (Palumbo et al. 2023).
The purpose of this paper is to propose a discriminatory test of general relativity focusing on a parametrised-spherically-symmetric spacetime model of M87*, as in Bauer et al. (2022); Kocherlakota & Rezzolla (2022), and to address the image degeneracy between the nature of the compact object and a large library of physically motivated synchrotron emission profiles for an equatorial disc. We test Keplerian, radially infalling and mixed velocity fields, with a special interest in the first secondary image of the flow, the $n=1$ photon ring. The main restrictions of this preliminary work are the adoptions of a non-rotating black hole and of a geometrically thin accretion disc. Nonetheless, the latter hypothesis may be justified very close to the event horizon both in the framework of classic thick disc accretion models in the plunging region, below the cusp (Abramowicz et al. 1978) and through more recent arguments based on general relativistic magnetohydrodynamic (GRMHD) simulations (Chael et al. 2021). Also, we only simulate instantaneous time-averaged images and we do not explore the associated interferometric visibility as measured by the EHT. These constrictive assumptions can be relaxed in subsequent work by making use of already existing examples for an axisymmetric parametrised spacetime (Konoplya et al. 2016), models for thick accretion flows (Vincent et al. 2022) and astrophysical fluctuations (Lee & Gammie 2021).
This paper is organized as follows. First, Section 2 reviews the definitions of the image components. Then, Sections 3 and 4 detail our models for the compact object and the accretion disc. Next, we describe our simulated images in Section 5 and we analyse some of their distinctive features in Section 6. Finally, we comment on the gravity-astrophysics degeneracy in Section 7 and offer concluding discussion in Section 8.
2 Schwarzschild image features and observables
Figure 1: Critical curve (dark blue), $n=1$ lensing band (light blue) and inner shadow (brown) of a Schwarzschild black hole on the plane of sky (upper panels) seen at different inclinations and their impact parameters as a function of the polar angle on the screen (lower panels).
In this paper, we consider static and spherically symmetric parametrised black hole metrics, that is, deformations of the vacuum Schwarzschild solution of general relativity. This section reviews the properties of null geodesics in this standard spacetime. However, the notions explored here can be and have been analysed in the spinning case, see Lupsasca et al. (2024) for a general introductory review on Kerr black hole imaging. Here, all the expressions are written in geometrised units ($G=c=1$), so $M$, the mass of the black hole, has the dimension of a length. We recall that the radial position of a Schwarzschild event horizon, in Schwarzschild-Droste coordinates $(t,r,\theta,\varphi)$, is $r_{\mathrm{H}}=2M$.
2.1 Null geodesics and photon shell
While the equations of geodesic motion in the Schwarzschild [^2] Their study in the Kerr metric was initiated by Carter (Carter 1968) and it was recently revisited and completed (Gralla & Lupsasca 2020b). spacetime were first derived by Schwarzschild (Schwarzschild 1916) and Droste (Droste 1917), the explicit examination of null geodesics governing the trajectory of massless particles like photons is, as far as we know, due to Flamm (Flamm 1916).
The four null geodesic equations can be obtained by means of the expressions translating the symmetries of the spacetime. The two Killing vectors $\boldsymbol{\partial}_{t}$ and $\boldsymbol{\partial}_{\varphi}$ associated with the stationarity and axisymmetry of the Schwarzschild spacetime result in the conserved energy $E$ and conserved angular momentum $L$:
$E=-\boldsymbol{\partial}_{t}\,\cdot\boldsymbol{p}=-p_{t}\,,\quad\quad L=\boldsymbol{\partial}_{\varphi}\,\cdot\boldsymbol{p}=p_{\varphi}\,,\quad\quad\lambda\coloneqq\frac{L}{E}\>,$
where $p^{\mu}=\mathrm{d}x^{\mu}/\mathrm{d}s$ are the components of the 4-momentum of the photon in Schwarzschild-Droste coordinates, with an affine parameter $s$ that increases along the geodesic towards the future. The energy-rescaled angular momentum $\lambda$, or equivalently the impact parameter $b\coloneqq|\lambda|$, is properly defined for null geodesics reaching an observer at asymptotic infinity, the ones of interest here, for which $E>0$, and it entirely characterises them (Wald 1984). Also, preserving the spherical symmetry confines the photon on a timelike-planar hypersurface and its mass $m=0$ represents the fourth integral of motion.
Excluding worldlines which terminate in the black hole region, there are no stable bound orbits for photons, since $r(s)$ is either a monotonic function or has a single radial turning point. Besides, photons orbit at fixed Schwarzschild-Droste radius $\tilde{r}$ if and only if, in Schwarzschild-Droste coordinates (Hilbert 1917)
$\tilde{r}=3M\quad\text{and}\quad\tilde{b}=3\sqrt{3}M\>.$ (2)
As mentioned above, these circular orbits are unstable: at the slightest perturbation, the photon initially at $\tilde{r}$ either falls inside the event horizon or it escapes to infinity. The region of the spacetime spanned by all the circular photon orbits is called the photon sphere [^3] For Kerr BHs we talk of spherical orbits and photon shell (Teo 2003).. A photon winds more and more times, clockwise or anticlockwise according to the sign of $\lambda$, as it approaches the photon sphere, where geodesics loop indefinitely.
2.2 Critical curve and lensing bands
The critical curve is the theoretical image-plane closed curve depicted by the impact points of null trajectories possessing the same constant of motion $\tilde{\lambda}$ as the bound photon orbit that they asymptotically approach. Simplistically, the critical curve, also referred to as ‘apparent boundary’ (Bardeen et al. 1973), is the projection of the photon sphere on the plane of sky and it delimits two regions according to the impact parameter $b$ of the geodesics. Reversing time’s arrow, one can imagine that the photon is launched from the screen of the observer, at $r_{\mathrm{em}}>\tilde{r}$, towards the black hole: inside the critical curve, that is, in the so-called ‘black hole shadow’ defined by $b<\tilde{b}$, null geodesics cross the horizon, while outside of it, when $b>\tilde{b}$, photon trajectories are not trapped. We stress that ‘black hole shadow’ is an unfortunate naming convention because it does not generally match with the effectively darker region on the image, as it is underlined in Paragraph 2.4. For a Schwarzschild black hole, the critical curve is a circle of radius equal to the critical impact parameter $\tilde{b}=3\sqrt{3}M\simeq 5.196152M$ [^4] In the Kerr case, its angular coordinates on the screen of a distant observer have been first derived by Cunningham and Bardeen (Cunningham & Bardeen 1973) and re-examined in present-day papers (Teo 2003; Johnson et al. 2020). Its cardioid shape, of similar size as the Schwarzschild one, depends on spin and inclination (Falcke et al. 1999).; we represent the critical curve as a dark blue line in Figure 1.
Lensing bands are regions on the observer’s screen that comprise all the impact points reached by photons with constants of motions close to those of a bound trajectory, that is, with $\lambda\approx\tilde{\lambda}$, and that follow nearly bound null orbits (Gralla et al. 2019): these geodesics pass close to the photon shell, without being trapped, in such a way that they can perform several turns around the black hole before leaving to infinity. More precisely, given an equatorial plane passing through the black hole, we say that a light ray belongs to the $n$-th lensing band, with $n\geq 1$, if and only if it intersects the given equatorial plane exactly $n+1$ times on its way to the observer (Chael et al. 2021; Paugnat et al. 2022). If the plane is parallel to the screen, this is equivalent to the statement that the photons of $n$-th lensing band execute $n$ half turns around the black hole before reaching the observer. See Appendix A for an alternative definition of lensing bands.
Lensing bands are perfectly circular for a black hole seen face-on, whereas they are highly deformed for increasing inclinations $i$ of the distant observer [^5] In a Kerr spacetime, they also weakly depend on the spin (Gralla & Lupsasca 2020a)., as shown by the light blue bands of Figure 1. In particular, lensing bands of an inclined screen possess a reflection symmetry [^6] This is only true for spherically symmetric spacetimes (Gralla & Lupsasca 2020a)., they correspond to those of the non-inclined screen along the line perpendicular to the axis of reflection and their width varies according to the polar angle on the screen; see Appendix A for a proof of these statements. Also, the inner edge of the $n$-th lensing band is traced by means of the impact parameters of light rays of order $n$ that asymptotically approach the event horizon, $r\to r_{\mathrm{H}}$, while its outer edge corresponds to the image of a circle of radius $r\to\infty$ on the plane transported by light rays of order $n$, see Figure A.3 in Appendix A. We highlight that in our plots, the outer edge of the lensing band is prescribed by large, but not infinite, values of $r$ according to numerical stop conditions on the disc extent.
The critical curve, as well as the lensing bands defined with respect to a given equatorial plane [^7] See Vincent et al. (2022) and Kocherlakota et al. (2024a) for a discussion on image orders when the crossed shape possesses some thickness. depend exclusively on the geometry of spacetime and are insensitive to the accretion disc properties. Nonetheless, because of the asymptotic character of their precise definitions, they are not the observables that could be directly measured from images of black holes.
2.3 Apparent shape of circular rings
Figure 2: Direct (solid lines in the upper panels) and first-lensed (dashed lines in the lower panels) apparent positions of rays emitted from isoradial distances $r$ to the black hole at different inclinations: $r=2M$ (grey), $2.5M$ (magenta), $3M$ (green) or $5.5M$ (yellow). The inner shadow (brown) and the first lensing band (light blue) are represented as well.
We analyse projections of equatorial rings of constant radii $r$. At this point we are only interested in mapping between a given equatorial plane and the observer’s screen, connected with null geodesics. Hence, we do not need to specify neither emission and opacity properties nor redshift effects. Furthermore, we restrain our attention to light rays of orders $n=0,1$, but the following enquiries can be generalised to higher order. Discussions presented herein are illustrated in Figure 2.
If the equatorial plane is observed face-on, the projections of a ring of radius $r$ within the disc are circles on the plane of sky. The arrival impact parameter $b$ of a photon emitted from a radius $r>r_{\mathrm{H}}$ only depends on $r$ and can be approximated by ‘just adding one’ (Gates et al. 2020). This relation, valid for a Kerr black hole, can be rewritten for a Schwarzschild black hole as
[…]
Hence, the event horizon, situated at the radius $r=2M$, will appear at an apparent approximated radius $b\approx 2.82M$, in agreement with the numerical outputs of Figure 2, while outer radii develop towards higher values of $b$. As for geodesics of order $n=1$, if one stops the backward ray-tracing integration, properly introduced in Section 5, before getting the third crossing of the plane, the numerical inner and outer edges of the $n=1$ lensing band will appear at $b=5.02M$ and $b=6.17M$ respectively (Gralla et al. 2019); these values agree with Figure 2.
If the screen and the equatorial plane are not parallel, the images of equatorial rings are deformed by projection effects [^8] In a Newtonian context, they would give rise to ellipses.. On the contrary, the critical curve does not vary under change of the plane’s inclination $i$ because of the spherical symmetry of the photon sphere and it always takes the shape of a circle of radius $b=3\sqrt{3}M$ on the plane of sky [^9] In Kerr, the photon shell has an equatorial reflection symmetry and the critical curve is always symmetric with respect to the axis perpendicular to the projected direction of the spin (Gralla & Lupsasca 2020a)..
Let us focus now on the general properties of projections on the observer’s screen of equatorial circles as seen by an inclined observer. The discussions of the following paragraphs are complemented by additional details given in Appendix A. First of all, both the zeroth-order and the first-lensed projections of planar circular rings are equal to their non-inclined counterparts along the line on the screen parallel to the inclination’s axis of rotation of the plane (horizontal axis in the upper panels of Figure 1). Moreover, projected rings always possess a mirror symmetry on the observer’s screen with respect to the axis that is normal to the inclination’s one [^10] This property, reliant on the planar motion typical of all static spherical symmetric spacetimes, is broken in Kerr (Gralla & Lupsasca 2020a). (vertical axis in Figure 1).
Let $\psi$ be the polar angle on the observer’s screen as shown in Figure 1. If the radius of the inclined ring satisfies $r_{\mathrm{H}}<r\leq\tilde{r}$, the zeroth-order projection of the side of the ring tilted backward ($0^{\circ}<\psi<180^{\circ}$, ) or forward ($180^{\circ}<\psi<360^{\circ}$) appears, respectively, at a higher or lower radial distances $b$ on the screen with respect to the non-inclined apparent circle. As a result, since in our simulations the forward side of the plane points towards $\psi=270^{\circ}$, the impact parameter $b$ of the inclined projected equatorial horizon, as well as that of rings with $r\leq 3M$, decreases from $\psi=90^{\circ}$ to $\psi=270^{\circ}$, see grey and magenta lines in the upper panels of Figure 2. The same behaviour of $b$ is followed for the forward part of rings with $r>3M$: it then decreases from $\psi=180^{\circ}$ to $\psi=270^{\circ}$ in our set-up, see green and yellow lines in the upper panels of Figure 2. Instead, for the backward points of rings with $r>3M$, for a given pair of plane’s inclination and direction on the observer’s screen, there exist a radius for which the behaviour of $b$ is reversed, that is to say that the projection of a given point has a smaller impact parameter than its corresponding non-inclined projection. The existence of this radius, which is smaller for higher inclinations and for directions on the screen closer to the axis of reflection, is a necessary consequence of the flat spacetime limit: in a Minkowski spacetime, the projected image of a circle is an ellipse whose semi-major axis is along the rotation axis (the horizontal direction here) and equal to the non-inclined projection. Since the semi-minor axis is smaller than the semi-major axis by definition, a reversal must appear at large distance from the black hole to reproduce the Minkowski behaviour. For the small inclination considered in the set-up of Figure 2, this effect appears beyond the considered field of view.
As for the first-lensed projection of equatorial rings, if the radius of the ring satisfies $r_{\mathrm{H}}<r\leq\tilde{r}$, the projection of the forward or backward side of the ring appears, respectively, at a higher or lower radial distances $b$ on the screen with respect to the non-inclined apparent circle. Hence, as the geodesic executes a half turn around the black hole before touching the screen and since in our simulations the forward side of the plane points towards $\psi=270^{\circ}$, the impact parameter $b$ of such rings decreases from $\psi=90^{\circ}$ to $\psi=270^{\circ}$, see grey and magenta lines in the lower panels of Figure 2. If $r=3M$, there is no variation with respect to the non-inclined case, see green line in the lower panels of Figure 2. Finally, if $r>3M$, the modulations of $b$ are reversed. Thus, in our simulations, $b$ increases from $\psi=90^{\circ}$ to $\psi=270^{\circ}$, see yellow line in the lower panels of Figure 2. Put differently, the ring projections flee the critical curve in the lower part of the screen and are attracted to it in the upper part.
2.4 Inner shadow, primary image and photon rings
The relevant observables appear on the far-away observer’s screen if and only if emitting material is present around the black hole. In particular, a single point source gives rise to an infinite sequence of images. These additional lensed observables were originally called ‘ghost images’ (Darwin 1959) and were present in the first simulated ‘photograph’ of an accretion disc around a Schwarzschild [^11] The first image of an accretion disc around a Kerr black hole with $a\neq 0$ has been computed by Viergutz (Viergutz 1993). black hole (Luminet 1979). SMBHs like Sgr A* and M87* are surrounded by emitting matter forming a hot, geometrically thick and optically thin radiatively inefficient accretion flow (RIAF, Yuan & Narayan 2014). Moreover, for M87*, EHT data favours the scenario of a magnetically arrested disc (MAD) whose millimetre wavelength emission arises near the event horizon, predominantly in the vicinity of the equatorial plane (Narayan et al. 2003; The Event Horizon Telescope Collaboration 2019c; The Event Horizon Telescope Collaboration 2021).
Images of black holes surrounded by optically thin emission display a brightness depression in their inner region, known as observable shadow, that does not always track the critical curve, but it is instead highly astrophysical dependent (Chael et al. 2021). For instance, for an idealistically thin equatorial disc, this ‘inner shadow’ can be well inside the critical curve: when the disc extends below the photon sphere, the edge of this darkest region corresponds to the projection of the inner radius of the equatorial disc modulated by redshift effects (Chael et al. 2021). For a disc seen face-on and extending to the event horizon of a Schwarzschild black hole, the inner shadow has a radius of about $2.82M$ on the observer’s screen, as discussed in Paragraph 2.3.
The primary image is the one reaching the observer on geodesics touching the emitting equatorial source just once. It is also called $n=0$ or direct image. Photon rings are higher order images of the same emitting material, labelled by the number $n\geq 1$ of the equatorial-plane crossings along their nearly bound orbit. When the source is intersected during the crossing, a flux increment is added. Also, the $n$-th photon ring of the equatorial source lies inside the $n$-th lensing band (Gralla & Lupsasca 2020a).
When the emitting source is not isotropically distributed, as in the case of equatorial discs, photon rings exponentially tend to the critical curve, have exponentially decreasing angular width and thus are exponentially demagnified and exponentially loose flux density [^12] The first M87* photon ring provides $\sim 10\%$ of the total luminosity in numerical GRMHD simulations (Johnson et al. 2020). for increasing $n$ (Johnson et al. 2020; Vincent et al. 2022). Also, in intensity cross sections of the images, they do not blend with the direct image, unlike an infalling sphere, but they give birth to a stack of discrete higher order images superimposed on top of the primary one (Gralla et al. 2019; Vincent et al. 2022), provided that absorption effects are weak at a given observing frequency (Beckwith & Done 2005). These properties are presented more meticulously in the introduction of Paugnat et al. (2022).
To summarise, photon rings are signatures related to the strong-field lensing effects of the photon sphere and, unlike the critical curve or the lensing bands, they can be observed and possess only limited astrophysical dependence (Vincent et al. 2022). The present EHT does not resolve the photon rings (Gralla 2021), but their interferometric detection may soon become possible either by observing at higher frequencies (ngEHT, Johnson et al. 2023) or on longer Earth-space baselines (BHEX, Johnson et al. 2024).
3 Compact object model
3.1 M87* distance and orientation
Our compact object model mimics the observationally assessed properties of M87*, namely its mass $M$, its distance $D$ and its inclination $i$. The values of $M$ and $D$ are those inferred by the EHT collaboration (The Event Horizon Telescope Collaboration 2019d), the mass being consistent with that obtained via a stellar-dynamics study (Gebhardt et al. 2011). For its inclination, defined as the angle between the line of sight and the angular momentum of the accretion flow, we use the value of the viewing angle of the jet in M87 (Mertens et al. 2016), assuming that its axis is aligned with the angular momentum vector of the disc and that the latter is pointing away from the observer.
3.2 Rezzolla-Zhidenko metric
In this paper, we consider the Rezzolla-Zhidenko metric (RZ metric, Rezzolla & Zhidenko 2014) which describes, in a parametric framework, the spacetime outside a general spherically symmetric [^13] The Konoplya-Rezzolla-Zhidenko metric (KRZ metric, Konoplya et al. 2016) generalises the RZ parametrisation to circular spacetimes, then to rotating black holes, but not all stationary axisymmetric black holes are included in this class. stationary, and therefore static, black hole [^14] For an extension of the RZ scheme to arbitrary asymptotically flat, spherical symmetric and static spacetimes, including non-BH ones like boson stars or naked singularities, see Kocherlakota & Rezzolla (2020).. In the strong-field regime, contrary to previous parametrisations in which an infinite number of equally important parameters is needed (Johannsen & Psaltis 2011), the RZ metric is described by a finite number of hierarchical factors. The line element of this metric in a spherical polar coordinate system $(t,r,\theta,\varphi)$ is:
$ds^{2}=-N^{2}(r)dt^{2}+\frac{B^{2}(r)}{N^{2}(r)}dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\varphi^{2})\>,$
where the functions $N$ and $B$ only depend on the radial coordinate $r$. Conveniently for a black hole, $r$ can be compactified in the dimensionless variable $x\coloneqq 1-r_{\mathrm{H}}/r$, where $r_{\mathrm{H}}$ is the radial position of the event horizon, so that $x=0$ corresponds to the location of the horizon and $x=1$ to spatial infinity. In terms of $x$, using $G=c=1$ natural units and introducing $\epsilon\coloneqq 2M/r_{\mathrm{H}}-1$, the functions $N$ and $B$ can be written as follows:
[…]
where $\tilde{A}$ and $\tilde{B}$ are approximated in the form of the following continued fractions:
$\tilde{A}(x)=\frac{a_{1}}{1+\dfrac{a_{2}\,x}{1+\dfrac{a_{3}\,x}{1+\ldots}}}\quad\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},\quad\tilde{B}(x)=\frac{b_{1}}{1+\dfrac{b_{2}\,x}{1+\dfrac{b_{3}\,x}{1+\ldots}}}\>\>\,.$ (6)
The dimensionless constants $a_{0},a_{1},a_{2},a_{3}\ldots$ and $b_{0},b_{1},b_{2},b_{3}\ldots$ have to be constrained by observations. If all $\epsilon$, $a_{i}$ and $b_{i}$ are zero, with $i\in\mathbb{N}$, the Schwarzschild metric is recovered in Schwarzschild-Droste coordinates.
It is useful to notice that at the horizon the functions $\tilde{A}$, $\tilde{B}$ of equations (6), reduce to $\tilde{A}(0)=a_{1}$, $\tilde{B}(0)=b_{1}$. This means that near the horizon, only the lower-order terms in the expansion are relevant. Higher-order terms increase the accuracy of the approximation and, for numerous known metrics deviating slightly from the Schwarzschild one, the first order is sufficient to accurately investigate astrophysical observable deviations (Konoplya & Zhidenko 2020). Besides, the coefficients $a_{0}$ and $b_{0}$, that have been introduced in the equations (5), can be expressed as a combination of the parametrised post-Newtonian (PPN) parameters and, assuming black hole uniqueness, they are constrained by observations to be at most of the order of $10^{-4}$ (Will 2014).
For the sake of simplicity, as we are interested in the near-horizon emission around a black hole and following the arguments stated before, in the simulations all the parameters $a_{i}$, $b_{i}$ with $i\geq 2$ are put to zero. The parameters $a_{0}$ and $b_{0}$ are also fixed to be zero, their impact on the image formation being negligible with respect to that of the astrophysical set-up when PPN constraints are met, as it is illustrated in Appendix C; in the same Appendix we also present higher deviations in $a_{0}$ and $b_{0}$, allowed in theories of gravity without a Birkhoff-like uniqueness theorem. Finally, we choose $b_{1}=0$, again for simplicity purposes and because its variation has a very limited effect on the image, as shown in Section 6 and Appendix C. Then, we can study the effect on the images of the only unconstrained parameters $a_{1}$ and $\epsilon$ that we will make vary independently, specifically fixing the other one to zero, in their theoretically-allowed [^15] The outermost horizon is located at $r=r_{\mathrm{H}}$ and no larger roots of $g_{rr}^{-1}=N^{2}(r)$ should exist (Kocherlakota & Rezzolla 2022). ranges and taking ‘small’, in the sense of the following right inequality [^16] This does not imply that the deviations are physically small., parametric deviations (Kocherlakota & Rezzolla 2020):
$-1<\epsilon\leq 0.5,\quad a_{1}\geq-1\quad\text{and}\quad\max{(\,|\,a_{1}\,|,|\,\epsilon\,|\,)}\leq 1\>.$
Table 1: Parameters of the black hole
Symbol Value Property
$M$ $6.5\times 10^{9}$ $\mathrm{M}_{\odot}$ Mass
$D$ 16.9 Mpc Distance
$i$ 163° Inclination
$a_{0}$, $b_{0}$, $b_{1}$, $a_{i}$, $b_{i}$ ($i\geq 2$) 0 RZ parameters
[…]
4 Accretion disc model
In this paper, the compact object is surrounded by an infinitely geometrically thin accretion disc orbiting in the equatorial plane taken to be at $\theta=\pi/2$ without any loss of generality. Our disc extends from an inner radius, $r_{\mathrm{inner}}$, given by the radial position of the event horizon, to a sufficiently large outer radius $r_{\mathrm{outer}}$ whose contribution to the image formation is negligible for our choice of quickly decreasing emission profiles.
4.1 Motion of the emitting material
We suppose that the emitters of mass $m_{\mathrm{e}}>0$ can have three possible types of dynamics: either the particles in the disc rotate around the black hole with a Keplerian velocity, they fall radially towards the central singularity or they possess a more realistic mixed velocity. A sketch of the two extreme velocity fields is reproduced in Figure 3. We define the specific energy $\varepsilon$ and specific angular momentum $\ell$ as
$\varepsilon\coloneqq\frac{E}{m_{\mathrm{e}}}\quad\text{and}\quad\ell\coloneqq\frac{L}{m_{\mathrm{e}}}\>.$
Figure 3: Illustration of the velocity fields and of the ray-tracing set-up. The orange massive particles of the yellow accretion disc can have a blue Keplerian velocity, be radially infalling along the red vectors or possessing a mixed velocity combining these two extreme cases. Schwarzschild-Droste coordinates are depicted in black, while the grey quantities locate the green screen of the observer.
4.1.1 Keplerian velocity
If the disc is Keplerian, the accreted matter follows a circular geodesic motion above the innermost stable circular orbit (ISCO), below which we assume that it spirals down along timelike geodesics with constants of motion $\varepsilon$ and $\ell$ equal to those of the ISCO (Cunningham 1975). A general timelike 4-velocity $u^{\,\mu}=p^{\,\mu}/m_{\mathrm{e}}=(u^{t},u^{r},0,u^{\varphi})$ can be expressed with the conserved quantities $\varepsilon$ and $\ell$ and via the normalisation $u^{\,\mu}u_{\mu}=-1$. Noting by primes the derivatives with respect to $r$, the expressions of the specific energy and z-component of the angular momentum for the RZ metric are then (Cárdenas-Avendaño et al. 2019)
$\varepsilon_{\mathrm{circ}}=\sqrt{\frac{N^{3}(r)}{N(r)-rN^{\prime}(r)}}\quad\text{and}\quad\ell_{\mathrm{circ}}=\sqrt{\frac{r^{3}N^{\prime}(r)}{N(r)-rN^{\prime}(r)}}\>,$
and the disc circular orbiters satisfy two additional conditions on the effective potential of the radial geodesic motion:
$V_{\mathrm{eff}}=0=V^{\prime}_{\mathrm{eff}}\quad\text{with}\quad V_{\mathrm{eff}}=\frac{\varepsilon_{\mathrm{circ}}^{2}}{N^{2}(r)}-\frac{\ell_{\mathrm{circ}}^{2}}{r^{2}}-1\>.$
The last stable orbit being marginally stable, $V^{\prime\prime}_{\mathrm{eff}}$, the second derivative of the effective potential, must vanish at the radius of the ISCO, $r_{\mathrm{ms}}$, which is then the real root of the following equation with no analytical solution (Rezzolla & Zhidenko 2014):
$3N(r_{\mathrm{ms}})N^{\prime}(r_{\mathrm{ms}})-3rN^{\prime 2}(r_{\mathrm{ms}})+rN(r_{\mathrm{ms}})N^{\prime\prime}(r_{\mathrm{ms}})=0\>.$
4.1.2 Radial infall
When the accreted matter is in radial free-fall towards the central compact object, the particle evolves in a constant-$\theta$ plane with vanishing specific angular momentum $\ell$. The conservation of the specific energy $\varepsilon$ and the normalisation of the 4-velocity lead to (Kocherlakota & Rezzolla 2022):
$u^{\mu}_{\mathrm{rad}}=\left(\frac{\varepsilon}{N^{2}},-\frac{\sqrt{\varepsilon^{2}-N^{2}}}{|\,B\,|},0,0\right)\>.$
We also neglect the radial velocity at infinity, so that:
$\lim_{r\to\infty}u^{r}_{\mathrm{rad}}=0\quad\text{and}\quad\varepsilon>0\quad\implies\quad\varepsilon=1\>.$
4.1.3 Mixed velocity
Realistic RIAF discs are neither totally radially infalling nor Keplerian, but, according to more sophisticated GRMHD simulations (Narayan et al. 2012), they likely consist of a mixed velocity comprising a radial component and a sub-Keplerian part (Yuan et al. 2022; Begelman et al. 2022). Moreover, the sub-Keplerian motion is expected to be predominant for the disc surrounding M87* (Chatterjee et al. 2023).
To define the sub-Keplerian velocity $u_{\mathrm{sub}}$ we follow the same reasoning of Paragraph 4.1.1, but we multiply the quantity $\ell/\varepsilon$ by the ‘sub-Keplerianity’ factor $\xi\in(0,1]$, so that above the ISCO
$\varepsilon_{\mathrm{sub}}=\sqrt{\frac{N^{3}(r)}{N(r)-\xi^{2}rN^{\prime}(r)}}\>,\quad\ell_{\mathrm{sub}}=\xi\sqrt{\frac{r^{3}N^{\prime}(r)}{N(r)-\xi^{2}rN^{\prime}(r)}}$
[…]
$u_{\mathrm{mix}}^{r}$ $=u_{\mathrm{sub}}^{r}+\left(1-\omega_{r}\right)\left(u_{\mathrm{rad}}^{r}-u_{\mathrm{sub}}^{r}\right)\>,$ (15.1)
$\Omega_{\mathrm{mix}}$ $=\Omega_{\mathrm{sub}}+\left(1-\omega_{\varphi}\right)\left(\Omega_{\mathrm{rad}}-\Omega_{\mathrm{sub}}\right)\>.$ (15.2)
Finally, $u^{t}_{\mathrm{mix}}$ is found by unit-normalisation and $u^{\varphi}_{\mathrm{mix}}=\Omega_{\mathrm{mix}}u^{t}_{\mathrm{mix}}$.
We note that this realistic mixed velocity does not correspond to a geodesic motion contrary to the purely Keplerian and radial ones, that are retrieved, respectively, by imposing $\omega_{r}=\omega_{\varphi}=1$ with $\xi=1$ and $\omega_{r}=\omega_{\varphi}=0$. In other words, the mixed velocity prescription can account for non-gravitational forces, that are expected to affect the emitters’ trajectory.
4.2 Synchrotron radiation
We assume the disc to be optically and geometrically thin, which means considering the optical depth to be idealistically null. The first hypothesis is only an approximation for the present EHT images of M87* and particularly for the photon rings observations, for which the opacity effects may play a role (Vincent et al. 2022). The assumption holds much better at 345 GHz, as absorption effects decrease with increasing frequency, and so it is reasonable for the observations at higher frequency of the ngEHT (Doeleman et al. 2023) or BHEX (Johnson et al. 2024). In the absence of absorption, the quantity $I_{\nu}/\nu^{3}$, with $I_{\nu}$ the intensity per unit frequency, is conserved along the geodesic (Lindquist 1966), so that the source specific intensity $I_{\nu}^{\mathrm{em}}$ is related to the observed one $I_{\nu}^{\mathrm{obs}}$:
$\frac{I_{\nu}^{\mathrm{obs}}}{I_{\nu}^{\mathrm{em}}}\propto\left(\frac{\nu^{\mathrm{obs}}}{\nu^{\mathrm{em}}}\right)^{3},$
where $\nu^{\mathrm{obs}/\mathrm{em}}$ is the observed or emitted frequency. Also, the quantity between parentheses represents the redshift factor:
$g\coloneqq\frac{\nu^{\mathrm{obs}}}{\nu^{\mathrm{em}}}=\frac{\boldsymbol{p}^{\mathrm{obs}}\cdot\boldsymbol{u}^{\mathrm{obs}}}{\boldsymbol{p}^{\mathrm{em}}\cdot\boldsymbol{u}^{\mathrm{em}}}\>,$
defined as the scalar product between the photon momentum at the observer or emitter, $\boldsymbol{p}^{\mathrm{obs}/\mathrm{em}}$, and the 4-velocity of the observer or emitter, $\boldsymbol{u}^{\mathrm{obs}/\mathrm{em}}$.
Even though plasma-physics phenomena, such as magnetic reconnection, can accelerate a population of electrons in power-law distributions, we assume here, to simplify, that the entire disc emits relativistic thermal synchrotron radiation; which is reasonable to characterise M87* emission at millimetre wavelengths [^18] The assumption is not appropriate for the infra-red flaring of Sgr A*..
Thus, according to equation (B3) in Vincent et al. (2022), and making the same approximations detailed therein in Appendix A, the relativistic electrons follow the Maxwell-Jüttner distribution at temperature $T_{\mathrm{e}}$, whose emissivity in cgs units is:
$j_{\nu}$ $\approx\frac{\sqrt{2}\pi e^{2}}{6c}\frac{n_{\mathrm{e}}\nu_{\mathrm{em}}}{\Theta_{\mathrm{e}}^{2}}\exp\left[-\left(\frac{9\pi m_{\mathrm{e}}c}{e\sin\theta_{\mathrm{B}}}\frac{\nu_{\mathrm{em}}}{B\Theta_{\mathrm{e}}^{2}}\right)^{1/3}\right],$ (18)
$\theta_{\mathrm{B}}$ $=\arccos(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})\>.$ (18.1)
where $n_{\mathrm{e}}$ is the number density of the electrons of mass $m_{\mathrm{e}}$, $e$ the charge of the electrons, $B$ the amplitude of the magnetic field, $\theta_{\mathrm{B}}$ the pitch angle between the unit vector along the direction of emission $\boldsymbol{\bar{K}}$ and the unit magnetic field vector $\boldsymbol{\bar{B}}$, both measured in the comoving frame of the emitting plasma [^19] See Appendix C.3 for a rigorous definition of $\boldsymbol{\bar{K}}$ and $\boldsymbol{\bar{B}}$., and $\Theta_{\mathrm{e}}\coloneqq k_{\mathrm{B}}T_{\mathrm{e}}/m_{\mathrm{e}}c^{2}$, $k_{B}$ being the Boltzmann constant.
Because the disc of our simulation is geometrically and optically thin, the radiative transfer equation reduces to:
$I_{\nu}^{\mathrm{em}}\propto j_{\nu}\>.$
In order to study the effects of the emission process on the simulated images, we assume that the physical quantities $n_{\mathrm{e}}$, $\Theta_{\mathrm{e}}$ and $B$ follow power-law radial distributions with indices $\alpha$, $\beta$ and $\gamma$ varying in ranges including the values obtained by 3D GRMHD simulations (Cho et al. 2023; Chatterjee et al. 2023) and in fair agreement with the ones set by Desire et al. (2025):
[…]
Substituting these power-laws in the emissivity, introducing the indices $i_{1}=\alpha-2\beta$ and $i_{2}=\gamma+2\beta$, gives:
$j_{\nu}$ $\approx\eta\frac{\nu_{\mathrm{em}}\text{\scriptsize[GHz]}}{230}\left(\frac{r}{r_{\mathrm{inner}}}\right)^{-i_{1}}\exp\left[-\zeta\sqrt[3]{\frac{\nu_{\mathrm{em}}\text{\scriptsize[GHz]}}{230\sin\theta_{\mathrm{B}}}}\left(\frac{r}{r_{\mathrm{inner}}}\right)^{i_{2}/3}\right],$ (21)
$\eta$ $\approx 1.31\times 10^{-18}\,\frac{n_{\mathrm{e};\mathrm{inner}}}{\Theta_{\mathrm{e;inner}}^{2}}\approx 10^{-3}\>\,\text{\footnotesize Jy}\>\text{\footnotesize cm}^{-1}\>,$ (21.1)
$\zeta$ $=\,\left(\frac{3.7\times 10^{5}}{B_{\mathrm{inner}}\Theta_{\mathrm{e;inner}}^{2}}\right)^{1/3}\in[2.5;4]\>.$ (21.2)
The value of $\zeta$, justified in detail in Appendix B of Vincent et al. (2022), is chosen in order to match the predicted characteristics of the accretion flow around M87*. The normalisation $\eta>0$, ensures that the 230 GHz radiative flux density is of the order of 0.5 Jy for a mean profile ($\zeta$=3, $\alpha$=1, $\beta$=2) as suggested by the EHT data (The Event Horizon Telescope Collaboration 2019b; Wielgus et al. 2020; The Event Horizon Telescope Collaboration 2024). Hence, our simulations only result in a correct flux density for this particular set of parameters. Nevertheless, since the normalization $\eta$ does not influence the conclusions of this work, we keep its value constant for simplicity.
Table 2: Parameters of the accretion disc
Symbol Value Property
$\xi$ 0.7 Sub-Keplerian parameter
$\omega_{\mathrm{r}}$, $\omega_{\mathrm{\varphi}}$ 0.8 Mixed-velocity parameters
$\zeta$2.5, 3, 4} Emission parameter
$\alpha$1, 1.5} Density power-law index
$\beta$1, 2} Temperature power-law index
$\gamma$ 1 Magnetic field power-law index
$\eta$ $10^{-3}$ Normalisation factor
$r_{\mathrm{inner}}$ $r_{H}$ Inner radius
$r_{\mathrm{outer}}$ 50 M Outer radius
 [^20] Fiducial parameters of a mean profile: $\zeta$=3, $\alpha$=1, $\beta$=2.
5 Adaptive backward ray-tracing
Figure 4: Adaptative ray-tracing polar grid with point densities adjusted for visual purposes. The real resolution values used in the simulated images are: angular resolution $\mathrm{res}^{\psi}=3.6^{\circ}$ (violet), $n=0$ radial resolution $\mathrm{res}^{b,\>n=0}=0.05$ µas (orange) outside the inner shadow (brown), $n=1$ radial resolution $\mathrm{res}^{b,\>n=1}=0.005$ µas (teal) in its lensing band (light blue). The star symbols represent the computed points of the apparent horizon and of the edges of the $n=1$ lensing band before interpolation. The inner 30 µas of a mean-profile sample images are also shown.
To simulate our images we use the ray-tracing code GYOTO (General RelativitY Orbit Tracer of the Observatoire de Paris, Vincent et al. 2011), collecting the specific intensity $I_{\nu}$ along null geodesics which are shot from pixels making up the screen of a distant observer towards the compact object. The resulting image is then a map of $I_{\nu}$ on the observer sky.
Keeping track of the equatorial crossings, it is possible to decompose the simulated image into different $n$ orders, properly defined for light rays in Section 2: the $n=0$ image is the map of the specific intensity obtained along the part of geodesic integrated before the second equatorial plane encounter, while the $n=1$ image is the map of $I_{\nu}$ accumulated between the second and the third crossing. We restrict our analyses to the zeroth and first orders so that the geodesic integration leading to the $n\geq 2$ images is not performed.
We define a polar coordinate system on the observer’s screen, that is, on the 2D simulated image, so that every point on it is described by its distance $b$ from the centre $O$ and by the positive angle $\psi$ from the horizontal axis passing from $O$, see Figure 4.
For each metric of Section 3, we produce adaptively ray-traced images based on the identification of the projection of the equatorial horizon and lensing bands introduced in Section 2, adapting the resolution for different image orders which are characterised by different scales. As explained in Gelles et al. (2021), this is an efficient way to explore the properties of computationally-expensive photon rings. To do that, we first simulate the image of a disc with constant emission and without spectral redshift effects ($\nu_{\mathrm{em}}=\nu_{\mathrm{obs}}$, $\zeta=0$ and $i_{1}=0$) on pixels separated by 0.1 µas along 14 equally spaced polar angular directions on sky. The disc is taken to have a very large effective outer radius in this case. Then, since the order of a null geodesic is related to the number of equatorial crossings, the structure of the image can be distinguished thanks to the number of impacts of the photon with the disc, itself linked to the value of $I_{\nu}$. In other words, for each angular direction, the boundary of the projected horizon is given by the further pixels from the centre having a null intensity, as no emission originates from the interior of the black hole by definition, and the edges of the $n=1$ lensing band are defined by the closest and furthest pixels cumulating the maximum intensity, since their geodesic collect twice the constant emission of the disc. Afterwards, the 14 computed values on the edges are connected with an interpolated curve. Next, for the actual emission profiles of Section 4, the $n=0$ image, computed for the pixels in the mask extending outside the apparent horizon until $r_{\mathrm{outer}}$ and the $n=1$ image, computed only inside the mask corresponding to the $n=1$ lensing band, with margins of 0.1 µas around the desired zone in order to be sure not to lose some needed pixels because of a bad initial resolution, are traced with adaptively chosen resolutions $\mathrm{res}^{b,\>n=0,1}$.
Our images are simulated at the two [^21] At 86 GHz absorption effects cannot be neglected. highest observing frequencies of the ngEHT and BHEX, 230 and 345 GHz (Doeleman et al. 2023; Johnson et al. 2024), and the screen orientation is chosen in such a way that the direction of the projected angular-momentum vector of the disc is aligned with the vertical axis. The field of view and the integration variables are chosen in order to have a reasonable compromise between a good image resolution and a modest computing time. The parameters of our simulations are given in Tables 1, 2 and 3.
Table 3: Parameters of the simulation
Symbol Value Property
$\nu^{\mathrm{obs}}$230, 345} GHz Observing frequency
$\mathrm{fov}$ 100 µas Field of view
$\mathrm{res}^{\psi}$ 3.6° Angular resolution
$\mathrm{res}^{b,\>n=0}$ 0.05 µas $n=0$ radial resolution
$\mathrm{res}^{b,\>n=1}$ 0.005 µas $n=1$ radial resolution
Figure 5: Dependence on the metric parameters $\epsilon$, $a_{1}$ of the event horizon, the photon sphere, the critical curve and the innermost stable circular orbit. The curves depending on $\epsilon$ or $a_{1}$ are traced at $a_{1}=0$ or $\epsilon=0$ respectively. Green lines indicate corresponding radii in a Schwarzschild spacetime.
6 Image analysis
Here, we consider an actual intensity map on the observer’s screen and we perform 100 1D intensity cuts equally spaced along the angular direction $\psi$. On each of them, with the same radial resolutions used to ray-trace the simulated primary image and $n=1$ photon ring, we compute the positions of the peaks and the widths of the profiles that are defined hereafter.
6.1 Position of the peak and width of an intensity profile
The peak along a given angular direction is the maximum of the 1D intensity profile. So, the location of the peak for a given $\psi$ is the impact parameter $b$ corresponding to the maximum intensity. Hereafter, we list the footprints on the intensity peak position of different astrophysical configurations or spacetime geometries, gradually incorporating complexities such as emission anisotropy and redshift effects.
If $\theta_{\mathrm{B}}$ is fixed to a constant value, the disc emission towards the observer is axisymmetric [^22] The local physics of the disc always remains axisymmetric.. Under this simplification, the emissivity $j_{\nu}$ of equation (21) and consequently the emitted intensity $I_{\nu}^{\mathrm{em}}$, related to $j_{\nu}$ via equation (4.2), reach their maximum at a radius $r_{\mathrm{max}}$ computed as
$r_{\mathrm{max}}=-\left(\frac{3i_{1}}{i_{2}\zeta}\sqrt[3]{\frac{230\sin\theta_{\mathrm{B}}}{\nu_{\mathrm{em}}\text{\scriptsize[GHz]}}}\right)^{3/i_{2}}r_{\mathrm{H}}\>.$
Given the range of parameters of Table 2, $r_{\mathrm{max}}\leq r_{\mathrm{H}}$ and $j_{\nu}$ is monotonically decreasing after $r_{\mathrm{max}}$. Therefore, all our emission profiles with $\theta_{\mathrm{B}}=\mathrm{const}$ radially decrease after $r_{\mathrm{H}}$. To put it another way, if we neglect frequency shift, that is, we put $g=1$ in equation (4.2), the observed intensity profiles $I_{\nu}^{\mathrm{obs}}$ are equal to the emitted ones, via equation (4.2), and peak, in this case, at the projected radial position of the event horizon for the $n=0$ emission and at the inner boundary of the $n=1$ lensing band, that is the $n=1$ image of the equatorial horizon as explained in Section 2, for the $n=1$ photon ring. Thus, without redshift effects, our constant-$\theta_{\mathrm{B}}$ observed peak positions can be predicted with the following ingredients:
  • The metric, which defines the event horizon, the photon shell and the geodesic motion, and consequently the related image features discussed in Section 2;
  • the inclination of the source, determining the deformations due to projection effects, see Section 2 and Appendix A.
For instance, in the case of a Schwarzschild black hole, the non-redshifted observed peak positions with $\theta_{\mathrm{B}}$ constant appear to be superposed on the solid, for $n=0$, or dashed, for $n=1$, gray lines in the plots of Figure 2 and they do not vary under change of the emission profile nor of the velocity of the disc.
Still in the non-redshifted framework, if the real value of the pitch angle $\theta_{\mathrm{B}}$, between the magnetic field vector and the emission direction in the rest frame of the emitter, is included in the computation of the emissivity $j_{\nu}$ of equation (21), the prediction of $r_{\mathrm{max}}$ is less straightforward and the axisymmetry of the emission profile is broken when the disc is not observed face-on; according to equation (18.1), the quantities needed to numerically compute $\theta_{\mathrm{B}}$ are then:
  • the emitted tangent 4-vector to the photon’s geodesic that depends both on the metric and the observer’s inclination (this results from lensing effects due to the spacetime curvature);
  • the direction of the magnetic field, taken to be everywhere vertical in our simulations. This assumption is defensible for a preliminary study of the compact emission from a MAD system, where the presence of a strong poloidal magnetic field is favoured (Narayan et al. 2003; The Event Horizon Telescope Collaboration 2021). Some results obtained considering the other component of a poloidal structure, namely, a radial magnetic field, are also shown in Appendix C.
If spectral effects are taken into account, the observed intensity $I_{\nu}^{\mathrm{obs}}$ is computed according to equations (4.2) and (4.2), so that it is both geometry-reliant and astrophysics-dependant:
  • the resulting $I_{\nu}^{\mathrm{obs}}$ is of course based on the shape of $I_{\nu}^{\mathrm{em}}$;
  • different 4-velocities of the emitter lead to different $g$ factors (this includes both the dynamical, or special relativistic, and the general relativistic redshift effects);
  • the emitted or observed 4-momentum of the photon depend on the metric and on the observer’s inclination (this is again the lensing effect due to the spacetime curvature).
As a result, the peak positions of our observed intensity profiles are shifted with respect to their corresponding non-redshifted impact parameters and this shift depends on the metric as well as on the properties of the accretion disc. Hence, already in the case of a Schwarzschild black hole with $\theta_{\mathrm{B}}$ fixed, the observed radii of the brightest parts of the images are no more superposed to the direct or lensed projections of the horizon, grey lines of Figure 2, and their values do vary under change of the emission profile and of the velocity of the disc.
We define the width of an intensity profile as its full width at half maximum, taking the difference between the impact parameters, above and below the peak’s one, whose associated intensity is closer to that of the maximum intensity divided by two.
The comments hereinabove regarding the dependencies of the redshift factor shaping the observed intensity $I_{\nu}^{\mathrm{obs}}$ are clearly still valid and relevant when considering its width, but contrary to the peak position which, at least for our class of synchrotron profiles and taking a constant $\theta_{\mathrm{B}}$, is invariant under change of the emission profile when spectral effects are neglected, the width is intrinsically dependent on the radiative properties of the disc.
6.2 Impact of the metric on the image features
We saw in Sections 2 and 6 that the horizon, the critical curve and the lensing bands are key elements to interpret black hole images, even though they are purely geometric, non-observable features. For instance, if we incorporate neither lensing effects in the emissivity nor frequency shift $g$, the peak positions of all our emission profiles track the apparent horizon for the $n=0$ image and the inner boundary of the first lensing band for the $n=1$ photon ring. In order to evaluate the geometrical effect on the image formation, it is then fundamental to understand how these characteristics vary when varying the spacetime metric.
We recall that the location of the event horizon under the RZ parametrization (Rezzolla & Zhidenko 2014) is only governed by the parameter $\epsilon$ and it is given by:
$r_{\mathrm{H}}=\frac{2M}{\epsilon+1}\>.$
Thus, $r_{\mathrm{H}}(\epsilon)$ is a positive decreasing function for $-1<\epsilon$ and the Schwarzschild event horizon, $r_{\mathrm{H}}=2M$, is retrieved for $\epsilon=0$, as it is shown in the first panel of Figure 5.
The radius $b_{\mathrm{c}}$ of the critical curve on the screen, or critical impact parameter, is well known for any spherically symmetric spacetime (Perlick & Tsupko 2022); for the RZ metric it reads:
$b_{\mathrm{c}}=\frac{r_{\mathrm{c}}}{|N(r_{\mathrm{c}})|}\quad\text{with}\quad\frac{\mathrm{d}}{\mathrm{d}r}\left(\frac{r^{2}}{N(r)^{2}}\right)\Bigg{\rvert}_{r_{\mathrm{c}}}=0\;.$ (24)
From equations (5), (6) we know that $N=N(r,\epsilon,a_{i})$, $i\in\mathbb{N}$, so $r_{\mathrm{c}}$ and $b_{\mathrm{c}}$ depend on these same parameters via equation (24). In particular, if we fix all the parameters to be zero except for $\epsilon$, $r_{\mathrm{c}}(\epsilon)$ and its projected radius $b_{\mathrm{c}}(\epsilon)$ decrease on the interval ]1;0.5]]-1;0.5]] - 1 ; 0.5 ], being $r_{\mathrm{c}}=3M$ and $b_{\mathrm{c}}=3\sqrt{3}M$ for $\epsilon=0$. If only $a_{1}$ varies, $r_{\mathrm{c}}(a_{1})$ and $b_{\mathrm{c}}(a_{1})$ are decreasing functions too and the Schwarzschild values are retrieved for a vanishing $a_{1}$. See the second and third panels of Figure 5 for a visual representation. We stress that the lensing bands always encompass the critical curve, so that its displacement sets in motion their inner edges.
Also the radial position $r_{\mathrm{ms}}$ of the innermost stable circular orbit only depends on $N(r,\epsilon,a_{i})$, $i\in\mathbb{N}$ via equation (4.1.1) and it plays a role in the computation of the 4-velocity profile of the emitter $\boldsymbol{u}^{\mathrm{em}}$ for a Keplerian disc, see paragraph 4.1.1, influencing then the value of its redshift factor $g$ via equation (4.2). On the treated intervals, both $r_{\mathrm{ms}}(\epsilon)$, and $r_{\mathrm{ms}}(a_{1})$, with vanishing $a_{1}$ or $\epsilon$ respectively, are decreasing functions, see last panel of Figure 5, and $r_{\mathrm{ms}}=6M$ for a Schwarzschild black hole with $\epsilon=a_{1}=0$.
To summarise, the values of $r_{\mathrm{H}}$, $b_{\mathrm{c}}$ and $r_{\mathrm{ms}}$ for a RZ non-standard black hole with positive individual parametric deviations, with $\epsilon\leq 0.5$, are smaller than in the Schwarzschild case, while they are bigger for $-1<\epsilon<0$, $-1<a_{1}<0$. This forces direct or lensed intensity’s peaks to follow the same tendencies.
Finally, we note that all the parameters of the RZ metric influence the null geodesic motion, in such a way that the impact locus on the screen of a trajectory reaching a given point around the black hole is modified by them. However, the parameters $b_{1}$ and $a_{0}$, $b_{0}$ (when PPN constraints are satisfied) engender minor modifications on the analysed ray-traced image characteristics, as it is reported in Appendix C.
6.3 Impact of the astrophysics on the observed intensity
Figure 6: Redshift profiles in a Schwarzschild spacetime for a disc with Keplerian velocity (solid lines) or radially infalling (dashed lines) seen face-on (left panel) or with inclination $i=163^{\circ}$(right panel). The $r$, $\varphi$ coordinates are those of the emission point of the photon. On the left panel, the emitted intensity with fiducial parameters and $\theta_{\mathrm{B}}=\mathrm{const}$ (dotted line) and its corresponding redshifted profiles (teal curves) are also traced.
We detail here the effects of various configurations of the disc on the observed intensity $I^{\mathrm{obs}}_{\nu}$ on which the measurements of the image analysis are performed, and on the key quantity needed to compute it, that is, the redshift factor $g$, see equations (4.2) and (4.2). We recall that when $g>1$ the observed frequency appears to be blueshifted, while if $g<1$, it is redshifted.
As mentioned above, at constant $\theta_{\mathrm{B}}$ all our emission profiles $I^{\mathrm{em}}_{\nu}$ are radially decreasing in the pertinent range, therefore peaking at $r_{\mathrm{inner}}=r_{\mathrm{H}}$. As for the observed profile $I^{\mathrm{obs}}_{\nu}$, given that the redshift factor $g$ always tends asymptotically to $g=1$, starting from $g=0$ at $r_{H}$, as it is recalled in Appendix B, the position of the peak of the observed intensity is moved to a further radius with respect to the one of the event horizon as it is shown in the left panel of Figure 6, reaching the screen with a corresponding larger impact parameter, see Figure 2. The observed shift is of course related to the viewing angle and the shape of the velocity profile of the disc. As for the width, if the gravitational redshift kills the profile on the left side of the peak, closer to the black hole, its right decay is mainly governed by the emission falloff.
If the disc is seen face-on, that is, if $i=0^{\circ}$, the redshift factor $g$ is just a function of the radial coordinate $r$, preserving the axisymmetry of the problem, and depends on the dynamics of the disc, see left panel of Figure 6: for every radial distance $r$, $g_{\text{Keplerian}}\geq g_{\text{Radial}}$ and both redshift factors tend to 1 from below, see left panel of Figure 6. Besides, photons along a given direction $\psi$ on the screen are emitted from a fixed $\varphi$ on the disc [^23] In Kerr, a given angle $\psi$ on the observer’s screen does not correspond to a specific angle around the black hole (Johnson et al. 2020)., see Figure 3. This means that the ray-traced intensity would be the same for every direction $\psi$ and, consequently, the intensity peaks are superposed to apparent isoradial curves like the ones illustrated in Figure 2 and the measured width would be constant along all the polar directions on the screen.
This azimuthal symmetry of the observed intensity is lost if the dynamical [^24] For a static disc, $g$ continues to be independent of $\varphi$, because in that case only the time components of the dot products are non-vanishing.disc is seen with some inclination and the $g(r,\varphi)$ profile of course still depends on the velocity of the disc, see right panel of Figure 6. Put differently, when some inclination is present, the well-known beaming effect, causing a boosting of the apparent luminosity of the part of the source moving towards the observer, emerges. It is visualised in Figure 3, the observer being located at $\varphi=270^{\circ}$: for the radial accretion flow velocity field, the approaching side corresponds to $\varphi\in[0^{\circ};180^{\circ}]$ (upper side of the image) with a maximum at $90^{\circ}$, whereas for a counterclockwise Keplerian velocity profile, the approaching side coincides with $\varphi\in[90^{\circ};270^{\circ}]$ (left side of the image), with a maximum at $180^{\circ}$. These variations, comparable to those of $\sin\varphi$ and $-\cos\varphi$ respectively, are coherent with the dependencies in $\varphi$ in the redshift factor when assuming a Minkowski spacetime, whose derivation is provided in Appendix B.
Figure 7: Positions of the $n=0$ (left panels) and $n=1$ (right panels) peaks, observed at 345 GHz and $i=163^{\circ}$, for all the possible emission configurations with fixed $\theta_{\mathrm{B}}=90^{\circ}$ in the case of a Schwarzschild black hole (green), a RZ black hole with $a_{1}|_{\,\epsilon=0}=0.1$ (pink) and a RZ black hole with $\epsilon\,|_{\,a_{1}=0}=0.1$ (violet). Each row corresponds to a different velocity: Keplerian, radial or mixed. Critical curves are represented as dashed lines.
7 Astrophysical-geometrical degeneracy
If one knew exactly the emission process and the dynamics that govern the disc, deductions on the underlying geometry could be undoubtedly made through the electromagnetic signatures of the accretion onto the black hole. Nonetheless, bearing in mind our ignorance of the complex astrophysical mechanisms involved (Gralla 2021), the question is to know if, in spite of all the astrophysics uncertainties, there are still some observables that permit to distinguish between a Schwarzschild spacetime and a non-standard metric. In the two following paragraphs, we explore the feasibility of the previous statement using the two measured quantities introduced in Section 6, namely the width and the peak position of the observed intensity along the polar angles on the observer’s screen. We additionally separate the simplistic case in which the angle $\theta_{\mathrm{B}}$, between the direction of emission and the magnetic field, is fixed to be $90^{\circ}$ from the one in which it is computed numerically via equation (18.1), in order to examine the importance of including the variation in the $\theta_{\mathrm{B}}$ angle caused by the spacetime curvature. The conclusions made are very similar for both the observing frequencies considered in Table 3, see Appendix C for an explanation, and the plots presented correspond to images simulated at $\nu^{\mathrm{obs}}=345$ GHz.
7.1 Observable disentanglement of the peak positions
The direct and first-lensed peaks positions for all possible combinations of the emission parameters of Table 2 are confined in the different bands of Figures 7 and 8, each band corresponding to one of the spacetimes considered in Table 3: Schwarzschild in green, RZ with $a_{1}=0.1$ and $\epsilon=0$ in pink, RZ with $\epsilon=0.1$ and $a_{1}=0$ in violet. The three investigated velocities of the disc, specifically Keplerian, radial and mixed with a sub-Keplerian predominant part plus radial component, are explored separately in the three rows of the plots.
Since the three bands are widely superposed in the first columns of Figures 7 and 8, we can state that astrophysical uncertainties dominate over small geometric deviations in the determination of the $n=0$ peaks and this regardless of the dynamics of the disc or the inclusion of $\theta_{\mathrm{B}}$. We remark that this is especially true for the Schwarzschild band, in green, and the $a_{1}|_{\,\epsilon=0}=0.1$ pink band which almost entirely cover each other because the event horizon, where all the emitted intensity profiles reach their maximum for a constant $\theta_{\mathrm{B}}$, see Section 6, is located at the same radial distance in these two cases. In other words, the other modifications induced by the non-standard metric, that is, on the geodesic motion, the redshift factor and the radius of the ISCO, see Section 6, are not sufficient to generate a significant observable difference on the $n=0$ peaks position and this is all the more true when varying only $b_{1}$, as illustrated in Appendix C. Even in the case of $\epsilon|_{a_{1}=0}=0.1$, whose event horizon’s radius is $r_{\mathrm{H}}\simeq 1.81M$, corresponding to a mean projected shift of about $0.84$ µas on the screen with respect to the Schwarzschild horizon, the measured positions of the peaks are highly degenerate between the green and violet bands. Higher parameters deviations lead, unsurprisingly, to more visible divergences, but even for the extreme positive values of equation (3.2), $a_{1}|_{\,\epsilon=0}=1$ and $\epsilon|_{a_{1}=0}=0.5$, the astrophysical-geometrical degeneracy is not broken for all the profiles, as shown in the first column of Figure C.7 of Appendix C.
Figure 8: Positions of the $n=0$ (left panels) and $n=1$ (right panels) peaks, observed at 345 GHz and $i=163^{\circ}$, for all the possible emission configurations computed including the real value of the synchrotron pitch angle $\theta_{\mathrm{B}}$ for a vertical magnetic field in the case of a Schwarzschild black hole (green), a RZ black hole with $a_{1}|_{\,\epsilon=0}=0.1$ (pink) and a RZ black hole with $\epsilon\,|_{\,a_{1}=0}=0.1$ (violet). Each row corresponds to a different velocity: Keplerian, radial or mixed. Critical curves are represented as dashed lines.
On the contrary, even being agnostic of the astrophysical emission, at least in the parameter space spanned in this paper, there is still the possibility to distinguish the strongly lensed observable features produced by different spacetime metrics. In other terms, as expected, the geometry of the spacetime impacts more dominantly the secondary images of the accretion disc. More precisely, for a non-standard black hole, the majority of the emission profiles with fixed $\theta_{\mathrm{B}}$ of a radially-infalling disc or of a disc with mixed velocity produce distinguishable peaks, see bottom right panels of Figure 7, and all of them become completely disentangled in the Keplerian case, see top right panel of Figure 7. Besides, if the real value of $\theta_{\mathrm{B}}$ is included, the degeneracy is always broken for all possible motions of the disc, as presented in the second column of Figure 8. This can be understood by the fact that subjacent spacetime plays an additional role in this second case, determining the lensing effects needed in the computation of $\theta_{\mathrm{B}}$. For a detailed explanation of the impact of $\theta_{\mathrm{B}}$ on the peak positions, also in the case of a radial magnetic field, see Appendix C. Again, as in the $n=0$ case, the distance from the Schwarzschild green band is more pronounced when the position of the event horizon is modified, as in the violet band [^25] We note that these violet peaks, contrary to the pink ones, would already be identified as a deviation from Schwarzschild when applying the purely geometric lensing band approach of Cárdenas-Avendaño & Held (2024), marginalising then over all conceivable emission profiles., and the contrasting conclusion vis-à-vis the $n=0$ study is mainly due to a modification of another key feature for the image formation, specifically the photon sphere around which $n=1$ light rays swirl before reaching infinity, see Section 6. This does not happen for deviations induced by the parameter $b_{1}$, see Appendix C. Here too, bigger parameters create more striking disentanglements, as shown in the second column of Figure C.7 of Appendix C. Also, a similar conclusion is reached when looking at higher order photon rings, see Figure C.1 in Appendix C.
As a remark, we note that no curve of Figure 2 is retrieved exactly because the redshift factor deforms the emitted profiles differently for various values of the azimuthal angle, see Section 6, so the measured peaks along the polar angle on the screen are not simply the projections of circular rings as discussed in 2.3. Finally, in Appendices C.4 and C.5, we complement our analysis by presenting the impacts of an inner radius of the disc that does not match the radial position of the event horizon or of a drastically different inclination of the observer on our conclusions.
7.2 Tangle of the widths
The width is not an observable allowing us to disentangle the effects of the astrophysics and the geometry on the electromagnetic images of accretion discs around supermassive black holes. For small deviations they are completely degenerate and even when taking extreme allowed values of the metric parameters, the degeneracy is globally preserved, this for every image order and disc velocity profile, see two last columns of Figure C.7. This is not startling in view of the properties already highlighted in Section 6. In fact, the width is measured by taking the difference between the two radial distances at which the intensity is half of the maximum: a smaller one, affected by the gravitational redshift, being indeed sensitive to the variation of the metric, and a larger one being mostly prescribed by the falloff of the emission profile. To a certain extent, it is fair to say that the parametric deviations of the spacetime are drowned in the diversity of the considered radiation parameters.
8 Discussion
8.1 Mass uncertainty
All the quantities of Figures 7 and 8 scale with the mass-to-distance ratio, $M/D$, which is expressed in angular units. From the data of the EHT 2017 campaign, the mass of M87* is measured to be $M=(6.5\pm 0.7)\times 10^{9}$ $\mathrm{M}_{\odot}$ under the assumption of GR and its distance is adopted to be $D=16.8\pm 0.8$ Mpc (The Event Horizon Telescope Collaboration 2019d). The precision on the mass reaches then about 11% whereas the one on the distance is close to 5%. Although the impact of the combined errors on $M$ and $D$ on our results should be examined, we note that the uncertainty on the mass is more than twice as important as that on the distance. Moreover, since the mass is the only parameter defining a Schwarzschild black hole, engendering the structure of its spacetime, we will assume, for the sake of simplicity, that the $D=16.9$ Mpc is perfectly known and that all the uncertainty on $M/D$ is concentrated on the mass. Thus, the question is to know whether a Schwarzschild black hole of lower mass, but still inside the EHT uncertainty range, could reproduce the $n=1$ characteristic peaks of a non-standard RZ black hole with the mass given in Table 1. To answer this, we put ourselves in the most optimistic scenario of Section 7, namely the one in which the disc has a mixed velocity and whose emissivity is given by expression (21), without further simplifications. Then, we superpose, in Figure 9, the bands of first-lensed peaks’ positions in the bottom right panel of Figure 8 to the ones produced by Schwarzschild black holes of lower masses: $M\in\{5.8,6.4,6.49\}$ $\mathrm{M}_{\odot}$, indicated by red ticks.
We see that the present EHT precision on the value of the black hole mass is not sufficient to discriminate the underlying metric [^26] The current $M/D$ uncertainty also disables metric tests with photon rings for all alternative spherically symmetric BHs in Wielgus (2021). because small RZ geometric deviations, in dotted bands, can be imitated by a standard black hole with lower mass, contained well inside the uncertainty range. In other words, according to the EHT results, the peaks’ band linked to a Schwarzschild spacetime can appear everywhere between the green bands. However, improved mass measurements with error bars of the order of $0.1\times 10^{9}$ $\mathrm{M}_{\odot}$, corresponding to 2% on the red colour bar, could not replicate the violet peaks produced by a RZ metric with $\epsilon|_{a_{1}=0}=0.1$ and all the small spacetime deviations considered in this paper could be detected with a precision on the black hole mass value of the order of $0.01\times 10^{9}$ $\mathrm{M}_{\odot}$. To resume, small geometric deviations modifying the position of the event horizon could be detected by increasing the precision on the measured mass by about one order of magnitude, whereas one order of magnitude more is needed to disclose a generic small spacetime deviation. This conclusion is unvaried when considering the $n=2$ photon ring, as displayed in Appendix C.
Conversely, assuming an infinite resolving power, the present EHT precision on the mass is sufficient to isolate the observables linked to the extreme $a_{1}|_{\epsilon=0}=1$ parametric deviation and the extreme positive deviation in $\epsilon$, $\epsilon|_{a_{1}=0}=0.5$, would become accessible too with an uncertainty on $M$ of about $0.4\times 10^{9}$ $\mathrm{M}_{\odot}$, that is, around 6% on the red colour bar. These results are illustrated by the starred bands in Figure 9. There exists then a one-to-one relationship between the value of the RZ parameters and the precision on $M/D$ needed to observe their effects.
Figure 9: Positions of the $n=1$ peaks, observed at 345 GHz, for all the possible emission configurations in the case of a black hole of mass $M=6.5\times 10^{9}$ $\mathrm{M}_{\odot}$ described by a Schwarzschild metric (dark green), a RZ metric with $a_{1}|_{\,\epsilon=0}=0.1$ or $1$ (dotted pink or starred pink) and a RZ metric with $\epsilon\,|_{\,a_{1}=0}=0.1$ or $0.5$ (dotted violet or starred violet), and a Schwarzschild black hole of mass $M=5.8\times 10^{9}$ $\mathrm{M}_{\odot}$ (light green). The velocity of the disc is taken to be mixed and the synchrotron pitch angle is computed in the emissivity. The red colour bar shows the error bar from the mean black hole mass value, $M=6.5\times 10^{9}$ $\mathrm{M}_{\odot}$, while the grey colour bars indicate the value of the RZ parameters $a_{1}$ (pink) and $\epsilon$ (violet).
8.2 Relative measurements
To go around the stringent precision needed for the $M/D$-dependent measurements of slightly deviating spacetimes discussed in Section 8.1, for an image-plane analysis, one has to consider relative measurements linked for instance to the image asymmetry or to respective size of subsequent images. In order to construct such quantities, we introduce the diameter $d^{n}$ of an $n$-order image as the distance between two intensity peaks along a polar direction on the screen. We then define the fractional asymmetry $A^{n}$ of an $n$-order image as one minus the ratio between its smallest and biggest diameters and the respective size $R^{n}$ of successive $n$ and $n+1$ images as the ratio between their mean diameters:
$A^{n}\coloneqq 1-\frac{d^{n}_{\mathrm{min}}}{d^{n}_{\mathrm{max}}}\quad\text{and}\quad R^{n}=\frac{d^{n+1}_{\mathrm{mean}}}{d^{n}_{\mathrm{mean}}}\>.$
The asymmetry of the critical curve has been already investigated in Johnson et al. (2020) and proven to not exceed a few percent for the spinning black hole in M87, being zero percent for a non-rotating central object. We find that, under the assumed inclination, the asymmetry is around 1% and 3% for the primary image and it reaches at most 0.3% for the first photon ring. The latter behaviour is well explained by the convergence of photon rings to the circular critical curve, exacerbated as expected for higher order images as described in Appendix C. We note that, both in the $n=0$ and $n=1$ asymmetries, the astrophysical uncertainties are strongly dominating, preventing any test of GR via this measurement, even for the extreme geometric deviations presented in Appendix C. However, if precisions of about 0.2% could be reached on the asymmetry of the primary image, some constraints could be put on the velocity distribution of the disc, reducing then the astrophysical assumptions of our models. Also, as the asymmetry of n=1 is highly dependent on the inclination of the source [^27] And to a lesser extent on spin, as the critical curve (Gralla & Lupsasca 2020a)., see Appendix C for more details, it could be used as a direct measurement of it, complementary to the present one where the disc is assumed perpendicular to the direction of the jet in M87 (Mertens et al. 2016).
Similar conclusions are reached for the size ratios of subsequent images: these quantities are highly degenerate between different spacetimes, and even more for high inclinations. One possibility to try to reduce this degeneracy is to consider a $M/D$-dependent prior on the emission radius (Wielgus et al. 2020). Anyway, although not directly useful to perform tests of GR, relative photon-ring sizes could be used to constrain the value of the spin of a rotating black hole (Broderick et al. 2022) and reduce then again the astrophysical uncertainty.
To summarise, it is unlikely to test RZ deviations from Schwarzschild with relative measurements. The main impediment is the fact that under the spherical symmetry assumption, RZ deviations seem to act essentially as scaling modifications. Therefore, we cannot exclude more promising outputs in more exotic spacetimes with, for example, deformed lensing bands or non-proportional shifts of the event horizon and the photon shell.
8.3 Conclusion
In this article, we have revisited the study already undertaken by Bauer et al. (2022) and Kocherlakota & Rezzolla (2022) on the prospect to test, via future interferometric observables of black holes like ngEHT (Johnson et al. 2023) or BHEX (Johnson et al. 2024), theories of gravity within the class of spherically symmetric RZ parametrised spacetimes (Rezzolla & Zhidenko 2014). Our main contributions to the effort is the inclusion of physically motivated synchrotron emission profiles in the equatorial plane (Vincent et al. 2022), parametrised according to the results of GRMHD simulations as in Desire et al. (2025), the simultaneous analysis of Keplerian, radial and mixed velocities of the thin accretion disc (Pu et al. 2016) and the simulations performed at the two higher observing frequencies of the ngEHT and BHEX, 230 and 345 GHz. We focus on the case of M87*, but our analysis could be generalised to other supermassive black-hole systems. Our findings can be summarised as follows:
  • all the measured features of the direct image are too polluted by the astrophysical uncertainty on the complex accretion process to be used as a robust probe of the spacetime geometry, consistently with the results obtained for spherical accretion models (Bauer et al. 2022; Kocherlakota & Rezzolla 2022), so that present EHT observations, even if performed at higher observing frequencies, do not have the necessary resolution to detect a violation of general relativity without additional constraints on the properties of the accretion disc. Stated differently, the discriminatory test proposed in this paper requires access to at least the $n=1$ photon ring;
  • the width of the $n=1$ intensity profiles along the polar angles on the screen are highly degenerate because the intrinsic emission profile of the disc governs their fall-off (Gralla et al. 2019); in other words, spacetime signatures on the width of the observed profiles can be mimicked by variations in the astrophysical configuration, so that the width of the $n=1$ photon ring is not a reliable quantity when looking for geometric modifications;
  • the peak positions of the $n=1$ intensity profiles along the polar angles on the screen are the only investigated features that allow to disentangle the deviations induced by the spacetime from those produced by a class of viable emission profiles [^28] Within GR, the $n=1$ photon-ring radii are mainly influenced by the Kerr BH spin, rather than by astrophysical diversity (Desire et al. 2025). Besides, our assertion holds for some, but not all, alternative spherically symmetric BHs in Eichhorn et al. (2023), $n=2$ being otherwise needed., in agreement to the conclusions of preliminary works (Kocherlakota & Rezzolla 2022). Also, the degeneracy is broken unambiguously, if the mass over distance is well constrained and for all possible motions of the disc, ranging from a Keplerian to a radially-infalling one, when the synchrotron emissivity is computed including the value of the synchrotron pitch angle $\theta_{\mathrm{B}}$, and this is true for both components of a poloidal magnetic field. Since the velocity of realistic accreted matter is expected to have a predominant Keplerian fraction, this result is promising, although challenging, for M87* spacetime’s constraints via future very-long-baseline interferometry (VLBI) missions;
  • Constraints on the black hole mass-to-distance estimate that are significantly more stringent than the present EHT ones, passing from an uncertainty on $M/D$ of about 11% (The Event Horizon Telescope Collaboration 2019d) to at least 2%, are needed to perform the discriminatory test of GR proposed in this paper using solely the intensity peak position of the first, or second, photon ring for small geometric deviations. On the other hand, some extreme metric deviations are already accessible with the current precision on $M/D$ given a sufficient resolving power;
  • Relative measurements, independent of $M/D$, such as image asymmetry or images size ratio, cannot be used directly to distinguish a spherically-symmetric conservative spacetime [^29] Relative measurements are sufficient for non-parametrised alternatives to the Schwarzschild metric (Wielgus 2021; Eichhorn et al. 2023).. Yet, they are complementary probes of the astrophysical set-up, which is at present the main impediment to the identification of small observational deviations from GR.
8.4 Limitations of this work and perspectives
A series of approximations are present in our modelling and the results presented only stand under such simplifying hypotheses.
First of all, our equatorial disc is taken to be optically and geometrically thin, so that absorption effects along the geodesic motion are neglected, and our simulations do not include any turbulence or observing noise. However, the luminosity of M87*, about $3.6\times 10^{-6}L_{\mathrm{Edd}}$ (Prieto et al. 2016), where $L_{\mathrm{Edd}}$ is the Eddington luminosity, is very low, corresponding to a radiatively inefficient accretion flow (Narayan et al. 1995), and RIAFs nuclei are expected to be surrounded by a hot, geometrically thick disc (Yuan & Narayan 2014). The integration of the above mentioned elements could on one hand weaken and spoil the lensed features of the image and on the other hand add astrophysical uncertainty and, thus, degeneracy. Nevertheless, EHT data favour the scenario in which the emission is dominated by that of the equatorial plane and at 345 GHz absorption effects are reasonably weak (The Event Horizon Telescope Collaboration 2019c), so that the framework of this paper remains adapted to model future VLBI observations. In any case, it is the objective of future work to expand the query of this paper to more realistic geometrically thick discs, computing the full radiative transfer (Vincent et al. 2022) and incorporating statistical fluctuations (Lee & Gammie 2021).
Secondly, we have only varied here one parameter at a time, fixing the others at their Schwarzschild value. Nonetheless, particular parametric combinations, taken of course into their theoretically-allowed ranges (Kocherlakota & Rezzolla 2022), may reduce the deviations induced by the metric, increasing then the entanglement with the astrophysical possibilities. If it would be worth analysing the outputs of such choices, it would be even more compelling to create a new framework in which specific parametric deviations have clear physical interpretations.
Third, the discussions made here concern image-plane time-averaged images, but it would be of great interest to know if the obtained results could be translated in terms of interferometric signals and be exploited for a direct real-data analysis, see for example Cárdenas-Avendaño & Lupsasca (2023).
Finally, we have only considered spherically symmetric spacetimes, but realistic black holes are expected to be rotating. Since the spin of M87* is not well constrained, this would represent an additional astrophysical parameter to scan and which may mimic the observational divergences produced by metric deviations. Nevertheless, the relative measurements presented in Section 8.2 could be used to measure its value (Johnson et al. 2020; Broderick et al. 2022). A generalisation of the RZ metric to circular axisymmetric spacetimes already exists (Konoplya et al. 2016) and its implementation in a ray-tracing code could be straightforward.
Acknowledgements.
I.U. and F.H.V. warmly thank Andrew Chael for insightful comments and constructive suggestions. M.W. is supported by a Ramón y Cajal grant RYC2023-042988-I from the Spanish Ministry of Science and Innovation.
Appendix A Inclination effects on projections
To understand lensing band deformations caused by the inclination between an equatorial plane and a distant observer, it is useful to employ the bending number $w$ introduced in Section 2.
Let us consider the set-up of Figure A.1. The far-away observer’s yellow screen, labelled by the polar coordinates $(b,\psi)$ as in Figure 1, is situated at $\varphi=\varphi_{\infty}=0\bmod 2\pi$ rad. We define a $\psi$-direction on the observer’s screen as a line passing through its centre with a slope defined by the polar angle $\psi$. From the same arguments exposed in Section 2, it is immediate to see that photons reaching the same $\psi$-direction on the screen, belong to the same orbital plane. Let us also consider an infinitely large green plane cutting the black hole in two hemispheres and whose approaching side to the observer is directed towards the south of his screen [^30] Different orientations of the plane can be easily deduced by a corresponding rotation of the north of the screen.. The approaching side of the plane, rotated around the east-west axis, is inclined at an angle $i<\pi/2$ rad [^31] If the equatorial plane contained a disc, the latter is not seen edge-on. from the face-on grey plane. We call $\Sigma$ the red angle between the intersections, with the orbital plane, of the inclined plane and of the face-on plane. For the east-west $\psi=0$ direction on the screen, $\Sigma$ is zero because the pink plane passes precisely along the intersection of the green and grey ones. For the north-south $\psi=\frac{\pi}{2}$ direction, $\Sigma$ is equal to $i$, the relative inclination between the green and grey planes. When rotating the orbital plane from east to north, $\Sigma$ takes intermediate values; the orbital plane of Figure A.1 corresponds indeed to an intermediate $\psi$-direction, for which $0<\psi<\pi$.
Figure 10: Intersections between a face-on (grey) or inclined (green) plane, the yellow screen of a distant observer and a pink orbital plane.
More precisely, the angle $\Sigma$ varies with the direction encountered by photons on the observer’s screen as follows:
$\cases{\>}\Sigma=0$ for ψ=0
Σ=i for ψ=π2
Σ=arccos{1+tanψ-21+tanψ-2+tan2i} otherwise. (A1)
So, $\Sigma$ increases symmetrically from $\psi=0$ to $\psi=\pi/2$ and from $\psi=\pi$ to $\psi=\pi/2$, as shown in Figure A.2.
Figure 11: Evolution of $\Sigma$ with the direction on the observer’s screen.
Figure 12: Null geodesics in a Schwarzschild spacetime reaching an observer on the far right. According to their winding number, geodesics can make up the inner shadow (brown), the $n=0$ image (orange) or the $n=1$ photon ring (light blue). We represent the projections of the black hole (black), the photon sphere (blue) and the equatorial plane for different polar directions on the screen: $\psi=0^{\circ}$ (pink), $45^{\circ}$ (red) and $90^{\circ}$ (violet).
Let us define the total change $\Delta\varphi$ of the orbital plane angle $\varphi$ and the bending number $w$ as in Gralla et al. (2019):
$\Delta\varphi\coloneqq\varphi_{\infty}-\varphi_{-\infty}\>,\>\>\>\varphi_{\pm\infty}=\lim_{\hat{s}\to\pm\infty}\varphi\quad\text{and}\quad w\coloneqq\Delta\varphi/2\pi\>.$ (A2)
Calling $\varsigma_{\mathrm{\lambda}}$ the sign of $\lambda$, we remark that $w=0$ for radial trajectories, $w\to\varsigma_{\mathrm{\lambda}}1/2$ when $b\to\infty$, while $w\to\varsigma_{\mathrm{\lambda}}\infty$ for $b\to\tilde{b}$. Having defined the bending number $w$ in equation (A2), we propose hereinbelow a classification of lensing bends which generalises previous conventions (Gralla et al. 2019; Chael et al. 2021) to arbitrary plane inclinations and polar directions on the observer’s screen. With $\varsigma_{\mathrm{\lambda}}$ the sign of $\lambda$ and $\sigma\coloneqq\Sigma/(2\pi)$, light rays are classified according to their order $n$, defined as follows: Thus, for $n\geq 1$, the $n$-th lensing band is the area on the plane of sky including all light rays whose bending number $w$ is inside the range settled by the order $n$, crossing then the equatorial plane exactly $n+1$ times along their way from the black hole to the observer or vice-versa, see Figure A.3. Given this definition, the periodicity of $\Sigma$, and so of $\sigma$, leads to the reflection symmetry of the lensing bands with respect to the north-south vertical direction on the observer’s screen. Put differently, if the projections of the equatorial plane in Figure A.3 tilt to the left from $\psi=0^{\circ}$ to $\psi=90^{\circ}$, they symmetrically come back to the initial perpendicular direction going towards $\psi=180^{\circ}$, imposing then the same geodesic orders.
Appendix B Redshift factor
B.1 Asymptotes
Let us discuss the limits of $g$ at infinity and at the event horizon. At infinite distance from the black hole, it is legitimate to consider that the emitter is static and that the spacetime is flat, in other words $\boldsymbol{u}^{\mathrm{em}}=\boldsymbol{u}^{\mathrm{obs}}$ and $\boldsymbol{p}^{\mathrm{em}}=\boldsymbol{p}^{\mathrm{obs}}$. By means of equation (4.2), the redshift factor at radial infinity is:
$\lim_{r\to\pm\infty}g=1\>.$
If the emitter is asymptotically close to the radius $r_{\mathrm{H}}$ of the event horizon, the radial component of its 4-velocity, $u^{r}_{\mathrm{em}}$, cannot be zero because below the radius of the ISCO, $r_{\mathrm{ms}}>r_{\mathrm{H}}$, no stable circular orbit is allowed and the 4-momentum of the emitted photon has to possess a radial component $p^{r}_{\mathrm{em}}$ to escape the gravitational attraction of the black hole and reach the observer at infinity. Besides, the 4-velocity of the emitter is proportional to its 4-momentum and 4-momenta’s time and azimuthal components are related to the constants of motion via equations (2.1). Then, at the horizon, the non-vanishing metric components of the RZ metric (3.2), (5), (6), with $x=1-r_{\mathrm{H}}/r$, read:
$\lim_{r\to r_{\mathrm{H}}}g_{tt}$ $\coloneqq\lim_{r\to r_{\mathrm{H}}}-N^{2}(r)=0\>,$ (B2.1)
$\lim_{r\to r_{\mathrm{H}}}g_{rr}$ $\coloneqq\lim_{r\to r_{\mathrm{H}}}B^{2}(r)/N^{2}(r)=\infty\>,$ (B2.2)
$\lim_{r\to r_{\mathrm{H}}}g_{\theta\theta}$ $\coloneqq\lim_{r\to r_{\mathrm{H}}}r^{2}=r_{\mathrm{H}}^{2}\>,$ (B2.3)
$\lim_{r\to r_{\mathrm{H}}}g_{\varphi\varphi}$ $\coloneqq\lim_{r\to r_{\mathrm{H}}}r^{2}\sin^{2}\theta=r_{\mathrm{H}}^{2}\sin^{2}\theta\>.$ (B2.4)
Thus, the scalar product $\boldsymbol{p}^{\mathrm{em}}\cdot\boldsymbol{u}^{\mathrm{em}}=p^{\mathrm{em}}_{\,\mu}\,u_{\mathrm{em}}^{\,\mu}$, appearing at the denominator of equation (4.2), diverges because the constants of motion are finite and one can show that so are $p_{r}^{\mathrm{em}}$ and $u^{r}_{\mathrm{em}}$, by considering, respectively, the normalisation $p^{r}_{\mathrm{em}}p_{r}^{\mathrm{em}}=0$ and the photon’s geodesic equation along $\theta$ or the expression (4.1.2) of a fully radial infall. Thus, when $r$ tends to the radial position of the event horizon, for every equatorial velocity profile of the emitter $\boldsymbol{u}^{\mathrm{em}}$, the limit of the redshift factor is:
$\lim_{r\to r_{\mathrm{H}}}g=0\>.$
B.2 Minkowski approximation
The Minkowski metric line element in spherical coordinates is:
$ds^{2}=-dt^{2}+dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\varphi^{2})\>.$ (B4)
We recall that the redshift factor $g$ is given by (4.2):
$g=\frac{\boldsymbol{p}^{\mathrm{obs}}\cdot\boldsymbol{u}^{\mathrm{obs}}}{\boldsymbol{p}^{\mathrm{em}}\cdot\boldsymbol{u}^{\mathrm{em}}}\>.$ (B5)
In Minkowski spacetime, the 4-momentum of the photon has a simple, geodesic invariant, analytical form (Vincent et al. 2024):
$\boldsymbol{p}^{\mathrm{obs}}=\boldsymbol{p}^{\mathrm{em}}=\mathbf{e}_{t}-\sin\iota\sin\varphi\,\mathbf{e}_{r}+\cos\iota\,\mathbf{e}_{\theta}-\sin\iota\cos\varphi\,\mathbf{e}_{\varphi}\>,$ (B6)
where $\iota=\pi-i\in[0^{\circ};90^{\circ}]$ for $i\in[90^{\circ};180^{\circ}]$ and where the natural basis vectors $\boldsymbol{\partial}_{\mu}$ have been replaced by:
$\mathbf{e}_{t}=\frac{\boldsymbol{\partial}_{t}}{\sqrt{-g_{tt}}},\quad\mathbf{e}_{r}=\frac{\boldsymbol{\partial}_{r}}{\sqrt{g_{rr}}},\quad\mathbf{e}_{\theta}=\frac{\boldsymbol{\partial}_{\theta}}{\sqrt{g_{\theta\theta}}},\quad\mathbf{e}_{\varphi}=\frac{\boldsymbol{\partial}_{\varphi}}{\sqrt{g_{\varphi\varphi}}}\>.$ (B7)
As the velocity of a static observer is directly given by $\boldsymbol{u}^{\mathrm{obs}}=\mathbf{e_{t}}$, one has $\boldsymbol{p}^{\mathrm{obs}}\cdot\boldsymbol{u}^{\mathrm{obs}}=-1$ at the numerator of equation (B5). As for the scalar product appearing at the denominator of $g$ in equation (B5), it varies according to the 4-velocity of the emitter. The general expressions of a Keplerian and a radial 4-velocities are:
$&\boldsymbol{u}_{\text{Keplerian}}=u^{t}_{\text{Keplerian}}(\boldsymbol{\partial}_{t}+\Omega\boldsymbol{\partial}_{\varphi})\>,\quad&&\Omega=u^{\varphi}/u^{t}=d\varphi/dt\\ &\boldsymbol{u}_{\text{Radial}}=u^{t}_{\text{Radial}}(\boldsymbol{\partial}_{t}+\chi\boldsymbol{\partial}_{r})\>,\quad\ &&\>\chi=u^{r}/u^{t}=dr/dt\>.\end{aligned}$ (B8)
The expressions of $\Omega$ and $\chi$ depend on the spacetime metric and $u^{t}$ can be deduced by the normalisation $\boldsymbol{u}\cdot\boldsymbol{u}=-1$. If one takes the expressions of $\Omega$ and $\chi$ for the Schwarzschild spacetime, which coincide with the Newtonian results, one gets:
$\Omega&=\frac{1}{r}\sqrt{\frac{M}{r}}&&\Rightarrow\quad u^{t}_{\text{Keplerian}}=\sqrt{\frac{r}{r-M}}\\ \chi&=-\sqrt{\frac{2M}{r}}&&\Rightarrow\quad u^{t}_{\text{Radial}}=\sqrt{\frac{r}{r-2M}}\>.\end{aligned}$ (B9)
Thus, in the orthonormal basis $\mathbf{e}_{\mu}$, the 4-velocities read:
$&\boldsymbol{u}_{\text{Keplerian}}=\sqrt{\frac{r}{r-M}}\left(\mathbf{e}_{t}+\sqrt{\frac{M}{r}}\mathbf{e}_{\varphi}\right)\\ &\boldsymbol{u}_{\text{Radial}}=\sqrt{\frac{r}{r-2M}}\left(\mathbf{e}_{t}-\sqrt{\frac{2M}{r}}\mathbf{e}_{r}\right)\>.\end{aligned}$ (B10)
With the previous derivations, it is then immediate to get the expression of the denominator of $g$
$&(\boldsymbol{p}^{\mathrm{em}}\cdot\boldsymbol{u}^{\mathrm{em}})_{\text{Keplerian}}=-\sqrt{\frac{r}{r-M}}\left(1+\sqrt{\frac{M}{r}}\sin\iota\cos\varphi\right)\\ &(\boldsymbol{p}^{\mathrm{em}}\cdot\boldsymbol{u}^{\mathrm{em}})_{\text{Radial}}=-\sqrt{\frac{r}{r-2M}}\left(1-\sqrt{\frac{2M}{r}}\sin\iota\sin\varphi\right)\end{aligned}$ (B11)
and so the analytical formula of the redshift:
$g^{\text{Minkowski}}_{\text{Keplerian}}&=\left[\sqrt{\frac{r}{r-M}}\left(1+\sqrt{\frac{M}{r}}\sin\iota\cos\varphi\right)\right]^{-1}\\ g^{\text{Minkowski}}_{\text{Radial}}&=\left[\sqrt{\frac{r}{r-2M}}\left(1-\sqrt{\frac{2M}{r}}\sin\iota\sin\varphi\right)\right]^{-1}\>.\end{aligned}$ (B12)
We note that for a disc seen face-on from below, $i=\pi\Rightarrow\iota=0^{\circ}$ and the Minkowski redshift factors are independent of $\varphi$.
Appendix C Additional parametric explorations
The aim of this section is to justify the choice of the relevant parameters considered in the main text and to survey a range of additional ones that go beyond the scope of this paper. All the simulations of this Appendix, unless differently stated, take into account the synchrotron pitch angle in the computation of the emissivity (21) and are performed at an observing frequency of 345 GHz. Also, we generally only consider the peaks position of the first photon ring of a disc with a mixed velocity profile around a black hole of mass $M=6.5\times 10^{9}$ $\mathrm{M}_{\odot}$ surrounded by a vertical magnetic field.
C.1 Observing frequency
The results of Figures 7 and 8, computed for $\nu_{\mathrm{em}}=345$ GHz, as well as the corresponding interpretations, are very similar to the ones obtained at 230 GHz, so that it is of no interest to plot or repeat them in the case of the higher observing frequency and we will just briefly explain why this resemblance is unsurprising. The observing frequency only intervenes in the computation of the observed intensity in equation (4.2), the emitted intensity and frequency being unchanged. Thus, value of the redshift factor of equation (4.2) is modified, but not its radial dependence, so that the main impact of the observing frequency is on the value of radiative flux density received in a given band and not on the peak position of normalised intensity cuts. Beyond GR tests, we recall the importance of observing at several frequencies to get physical insights on the complex and multi-wavelength emission process and to improve the interferometric coverage for image reconstruction (Johnson et al. 2023).
C.2 Higher order photon rings
Figure 13: Positions of the $n=2$ peaks for all the possible emission configurations in the case of a Schwarzschild black hole (green), a RZ black hole with $a_{1}|_{\,\epsilon=0}=1.0$ (pink), a RZ black hole with $\epsilon\,|_{\,a_{1}=0}=0.5$ (violet) and Schwarzschild black holes of mass $M=6.4\times 10^{9}$ $\mathrm{M}_{\odot}$ (hatched turquoise) and $M=6.3\times 10^{9}$ $\mathrm{M}_{\odot}$ (hatched teal). Critical curves are represented as dashed lines.
We simulate the second photon ring adaptively, in the $n=2$ lensing band, with a radial resolution $\mathrm{res}^{b,\>n=2}=0.001$ µas. This higher lensed observable appear to have more clearly separated intensity peaks positions given a sufficient resolving power. Nonetheless, as for the $n=1$ image, these peaks cannot be used as a discriminating probe of any conservative spacetime when the uncertainty on the mass-to-distance estimate is too high; here a precision of 3% is needed. This limited relaxation of the $M/D$ precision constraint can be understood by recalling that photon rings track the critical curve for increasing $n$ and that the position of the critical curve is a $M/D$ highly dependent quantity. The relative measurements introduced in Section 8.2 are again very degenerate even for this more strongly lensed observable, the asymmetry being of the order of 0.01% for all considered spacetimes.
C.3 Magnetic field configuration
As already mentioned in the main text, for a MAD system the presence of a strong poloidal magnetic field is favoured (Narayan et al. 2003; The Event Horizon Telescope Collaboration 2021). Thus, it is of interest to consider the radial component of the magnetic field in addition to the vertical one. In principle, the advection of the rotating fluid may also create a magnetic-field azimuthal component (Yuan & Narayan 2014), but the latter is not investigated in this paper.
Figure 14: Positions of the $n=1$ peaks for all the possible emission configurations in the case of a Schwarzschild BH (green), a RZ BH with $a_{1}|_{\,\epsilon=0}=0.1$ (pink) and a RZ BH with $\epsilon\,|_{\,a_{1}=0}=0.1$ (violet). Critical curves are represented as dashed lines and the magnetic field is radial.
Figure 15: $n=1$ geodesics emitted at $r=2,3,6,15\,M$ and directions of emission (blue) as well as those of a vertical (green) and a radial (red) magnetic field. The inset shows the value of the the pitch angle $\theta_{\mathrm{B}}$ of the first lensed photons for the two, vertical and radial, magnetic field configurations in the case of a disc with mixed velocity profile.
The peak positions obtained for a radial magnetic field, see Figure C.2, are equivalent in terms of astrophysical-geometrical degeneracy with respect to the ones assuming a vertical magnetic field in Figure 9. However, contrary to the vertical case, all bands in Figure C.2 are located strictly above the critical curve. Let us analyse the reasons of this different behaviour. First of all, it is worth noticing that the synchrotron pitch angle $\theta_{\mathrm{B}}$ only appears in the emissivity of equation (21) which is then the only quantity affected by a change of the magnetic field configuration. Also, from its definition, that is, the angle between the magnetic field vector and the emission direction in the rest frame of the emitter, it is clear that its value would be different for a same point source if one considers its direct or secondary image, because the directions of emission of a $n=0$ or $n=1$ photon differ. Let us then treat $n=0$ and $n=1$ separately and consider a disc seen face-on around a Schwarzschild black hole for simplicity [^32] We expect our statements to be just slightly modified when adding a small inclination or small metric deviations.. Let $\boldsymbol{k}=\boldsymbol{p}^{\mathrm{em}}/\hbar$, with $\hbar$ the reduced Planck constant, be the 4-vector tangent to the photon’s geodesic at emission. Its space-like projection $\boldsymbol{K}$ orthogonal to the 4-velocity $\boldsymbol{u}$ of the emitter, that is, in the rest-frame of the emitter, reads (Vincent et al. 2024): $\boldsymbol{K}=\boldsymbol{k}+(\boldsymbol{k}\cdot\boldsymbol{u})\,\boldsymbol{u}$. In the rest frame of the emitter, the magnetic field 4-vector is also orthogonal to the 4-velocity of the emitter: $\boldsymbol{B}\cdot\boldsymbol{u}=0$, so that $\boldsymbol{K}\cdot\boldsymbol{B}=\boldsymbol{k}\cdot\boldsymbol{B}$. Let us now define $\boldsymbol{\bar{K}}$ and $\boldsymbol{\bar{B}}$ as the unit vectors along $\boldsymbol{K}$ and $\boldsymbol{B}$ respectively. As $\boldsymbol{k}\cdot\boldsymbol{u}=-\omega$, with $\omega$ the positive pulsation of the photon as measured by the emitter, $\boldsymbol{\bar{K}}=\boldsymbol{K}/\omega$. As for $\boldsymbol{\bar{B}}$, considering that $\boldsymbol{u}=u^{t}(\boldsymbol{\partial}_{t}+\chi\boldsymbol{\partial}_{r}+\Omega\boldsymbol{\partial}_{\varphi})$, with $\chi=u^{r}/u^{t}$ and $\Omega=u^{\varphi}/u^{t}$, via equation (B7), it is straightforward to get [^33] We note that for a pure Keplerian velocity, $\chi=0$ and so $\boldsymbol{\bar{B}}_{\text{Radial}}=\mathbf{e}_{r}$.:
$&\boldsymbol{\bar{B}}_{\text{Vertical}}=-\mathbf{e}_{\theta}\\ &\boldsymbol{\bar{B}}_{\text{Radial}}=\frac{1}{\sqrt{-(g_{tt}+\chi^{2}g_{rr})}}(\,\chi\sqrt{g_{rr}}\,\mathbf{e}_{t}+\sqrt{-g_{tt}}\,\mathbf{e}_{r})\>.\end{aligned}$ (C1)
It follows then
$&(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})_{\text{Vertical}}=-\frac{k^{\theta}}{\omega}\sqrt{g_{\theta\theta}}\\ &(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})_{\text{Radial}}=\left(\,\frac{k^{r}}{\omega}-\chi\,\frac{k^{t}}{\omega}\right)\frac{\sqrt{-g_{tt}g_{rr}}}{\sqrt{-(g_{tt}+\chi^{2}g_{rr})}}\>.\end{aligned}$ (C2)
The synchrotron pitch angle is defined from these quantities as:
$\theta_{\mathrm{B}}$ $=\arccos(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})\>.$ (C3)
At very large radial distance on the disc, the emitter can be approximated as static, so that $\chi=0$ and $\omega=-\boldsymbol{k}\cdot\boldsymbol{u}=-g_{tt}k^{t}$. Also, since far from the black hole the spacetime is described by the Minkowski metric, with $g_{tt}=-1$ and $g_{rr}=1$, we get $\omega=k^{t}$ and $k^{t}$ can be related to the photon energy defined in equation (2.1): $k^{t}=g^{tt}k_{t}=-E/(g_{tt}\hbar)=E/\hbar$. Using all these remarks, equations (C2) drastically simplify in the asymptotic limit $r\to\infty$: $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{\infty}_{\text{Vertical}}=-\hbar rk^{\theta}/E$ and $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{\infty}_{\text{Radial}}=\hbar k^{r}/E$. Then, noticing that for a face-on observer the angular momentum $L$ of equation (2.1) is zero for all photons, one can retrieve the expressions of $k^{\theta}$ and $k^{r}$ by taking the vanishing spin parameter $a$ in the Kerr equations of motion (Gralla & Lupsasca 2020b). We finally get: $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{\infty}_{\text{Vertical}}=\mp_{\theta}\sqrt{\bar{Q}}/r$, with the minus sign corresponding to a photon emitted downwards in $\theta$, and $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{\infty}_{\text{Radial}}=\pm_{r}\sqrt{1-\bar{Q}(r^{2}-2Mr)/r^{4}}$, with the minus sign corresponding to a photon emitted towards the black hole radially. For an on-axis observer, the reduced Carter constant $\bar{Q}$ is directly related to the polar radius $b$ on the screen (Gralla & Lupsasca 2020a) via $\bar{Q}=b^{2}$ and the approximation (2.3), $b\approx r+M$, then yields: $\bar{Q}\approx r^{2}+2Mr+M^{2}$. Thus, for $n=0$ photons, $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{\infty,\,n=0}_{\text{Vertical}}=-1\Rightarrow(\sin\theta_{\mathrm{B}})^{\infty,\,n=0}_{\text{Vertical}}=0$ and $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{\infty,\,n=0}_{\text{Radial}}=0\Rightarrow(\sin\theta_{\mathrm{B}})^{\infty,\,n=0}_{\text{Radial}}=1$. For $n=1$ photons, it is clear from Figure C.3 that the radial growth of $\bar{Q}$, unchanged along one geodesic, is much slower than $r^{2}$ with respect to the second impact radius. As a consequence, $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{\infty,\,n=1}_{\text{Vertical}}=0\Rightarrow(\sin\theta_{\mathrm{B}})^{\infty,\,n=1}_{\text{Vertical}}=1$ and $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{\infty,\,n=1}_{\text{Radial}}=-1\Rightarrow(\sin\theta_{\mathrm{B}})^{\infty,\,n=1}_{\text{Radial}}=0$, where the ingoing sign can be understood from Figure C.3. When radially approaching the event horizon, no trivial [^34] Except for the conditions $u^{\theta}=0$ and $L=p_{\varphi}=g_{\varphi\varphi}p^{\varphi}=0$, translating an equatorial disc motion and a face-on view respectively. simplifications apply to $\omega=-g_{tt}k^{t}u^{t}-g_{rr}k^{r}u^{r}$. However, again from the geodesic equations (Gralla & Lupsasca 2020b), the asymptotic behaviour of $w$ is $w^{r_{\mathrm{H}}}=[u^{t}-u^{r}r/(r-2M)]E/\hbar$. Then, using the fact that $u^{t}>0$ for future-directed particles and $u^{r}$ is negative and non-vanishing below the radius of the ISCO, it is straightforward to conclude that $\omega\to+\infty$ at the horizon. As $k^{\theta}$ is finite at the horizon, for a vertical magnetic field, we find: $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{r_{\mathrm{H}},\,n=0,1}_{\text{Vertical}}=0\Rightarrow(\sin\theta_{\mathrm{B}})^{r_{\mathrm{H}},\,n=0,1}_{\text{Vertical}}=1$. In the case of a radial magnetic field, we show that the asymptotic limit of equation (C2) is $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{r_{\mathrm{H}},\,n=0,1}_{\text{Radial}}=\varepsilon_{\mathrm{sub}}r\left(\varepsilon_{\mathrm{sub}}^{2}r^{2}+C^{2}\omega_{\varphi}^{2}\ell_{\mathrm{sub}}^{2}\right)^{-1/2}$ where we have introduced $C\coloneqq-\omega_{r}\varepsilon_{\mathrm{sub}}-(1-\omega_{r})$ and employed the quantities $\varepsilon_{\mathrm{sub}},\ell_{\mathrm{sub}},\omega_{r}$ and $\omega_{\varphi}$ presented in Paragraph 4.1.3. For a sub-Keplerian motion, for which $\omega_{r}=\omega_{\varphi}=1$, this quantity reduces to $r\left(r^{2}+\ell_{\mathrm{sub}}^{2}\right)^{-1/2}=\frac{1}{2}[(4-\xi^{2})/(1+2\xi^{2})]^{1/2}$ which is equal to 0.5 for Keplerian velocity profile with $\xi=1$, see Paragraph 4.1.3. For a radial infall, the previous scalar product becomes 1 as $\omega_{r}=\omega_{\varphi}=0$. To summarise, given the values of Table 2, $(\sin\theta_{\mathrm{B}})^{r_{\mathrm{H}},\,n=0,1}_{\text{Radial}}\simeq 0.87$, 0 or 0.67 for a Keplerian, radial or mixed velocity profile respectively. Let us finally consider the intermediate extrema of the $\sin\theta_{\mathrm{B}}$ function. Generally, an extremum of $\sin\theta_{\mathrm{B}}$ is found either when $\cos\theta_{\mathrm{B}}=0\Leftrightarrow\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}}=0$ or when $\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}}\neq\pm 1$ and $d(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})/dr=0$. The first condition, corresponding to a maximum as $\sin\theta_{\mathrm{B}}=1$, is never verified for a vertical magnetic field, as $\omega$ is finite and $k^{\theta}\neq 0$ for non-extremal values of $r$, and it is satisfied, in the case of a radial magnetic field for $k^{r}/k^{t}=\chi=u^{r}/u^{t}$. We numerically show that, for all the velocity profiles investigated in this paper, this relation is never verified for the $n=0$ photons and only for a single value of $r$, depending on the disc velocity, in the case of $n=1$ photons. In particular, $r\simeq 3.6$, 5.6 or $4.4M$ in the case, respectively, of a Keplerian, radial or mixed velocity with the parameters of Table 2. In the case of a vertical magnetic field, using Leibniz rule and the geodesic equations (Gralla & Lupsasca 2020b), the second condition can be rewritten as $r=-\omega/(d\omega/dr)$. We note that this condition is never verified for $n=0$ photons, whereas it leads to a single [^35] As $(\sin\theta_{\mathrm{B}})^{n=1}_{\text{Vertical}}$ is not a constant function and it tends to 1 both at infinity and at the horizon, we already knew that it possesses at least one minimum; here, we claim the unicity of this extremum. minimum position $r>2M$ for $n=1$ photons: $r\simeq 3.9$, 5.3 or $4.6M$ for a Keplerian, radial or mixed velocity, respectively. For a radial magnetic field, no trivial simplification of the second condition can be easily found, but we notice that an additional minimum is present for the Keplerian and mixed velocity profiles both for the $n=0$ and $n=1$ photons at, respectively, $r\simeq 2.6M$ and $r\simeq 2.7M$ or $r\simeq 2.2M$ and $r\simeq 2.3M$. Let us now summarise our findings and their implications for the position of the peaks of intensity. For weakly lensed photons and a radial magnetic field, $\sin\theta_{\mathrm{B}}$ increases from an intermediate value to 1 with increasing distance from the central object, after reaching a minimum in the case of a Keplerian or mixed velocity, while it decreases from 1 to 0 in the presence of a vertical magnetic field. As a consequence, $j_{\nu}$ given by equation (21), at a fixed $r$, is smaller, in both cases, than when taking $\theta_{\mathrm{B}}=90^{\circ}$, but more importantly, the emission profile dies faster for a radial magnetic field and even more so for a vertical one. This implies that, when incorporating redshift effects, the $n=0$ peak positions will be reached at further radii for the $\theta_{\mathrm{B}}=90^{\circ}$ profiles than for the ones where $\theta_{\mathrm{B}}$ is computed numerically. Specifically, the maxima occur at smaller radii in the case of a radial magnetic field and even smaller radii for a vertical field. Part of this reasoning is confirmed by comparing the left columns of Figures 7 and 8. Concerning the first-lensed photons, whose geodesics are illustrated in light blue in Figure C.3, a global extremum in $\sin\theta_{\mathrm{B}}$ is present in all scenarios, being a minimum for a vertical magnetic field and a maximum for a radial one, see right inset of Figure C.3 for a mixed velocity profile. Then, in the radial-$B$ case, inner emission below the maximum is weakened even before the redshift knock down detailed in Section 6. This explains the behaviour remarked in Figure C.2: $n=1$ peak positions appear further away for a radial magnetic field. Similarly, as the inner emission for a vertical magnetic field is boosted, $n=1$ peaks appear at smaller radii than the ones of profiles with constant $\theta_{\mathrm{B}}=90^{\circ}$; see right columns of Figures 7 and 8. The integration of $\theta_{\mathrm{B}}$ in the thermal synchrotron emission profile is then an essential ingredient to take into account when interpreting observed images and its imprints on lensed observables could provide some hints on the magnetic field structure.
C.4 $r_{\mathrm{inner}}$
As already pointed out in Section 7.1, the key ingredient to break the degeneracy with the peak positions of the first photon ring is the different size of the critical curve in various spacetimes. The question now is to know how robust this result is when varying the position of the inner radius of the disc. For example, by taking a RZ spacetime with $\epsilon\,|_{\,a_{1}=0}=0.1$ and a disc with an inner radius $r_{\mathrm{inner}}=2M$ instead of $1.81M$, coinciding with the horizon of a Schwarzschild black hole, the violet band of the bottom right panel of Figure 8 is shifted upwards by at most 0.05 µas, preserving the conclusions about possible observable disentanglements. However, already for $r_{\mathrm{inner}}=3M$, the violet band is just 0.1 µas below the pink one, meaning that the impacts of $a_{1}$ and $\epsilon$ become comparable. The position of the inner radius of the disc is even more crucial for the analysis of the primary image: unlike variations in the critical curve, which do not impact the $n=0$ peak positions, the latter are primarily governed by the inner radius, where the considered emission profiles reach their maximum intensity. For instance, in the case, again, of a RZ spacetime with $\epsilon\,|_{\,a_{1}=0}=0.1$ and a disc with $r_{\mathrm{inner}}=2M$, the the violet band of the bottom left panel of Figure 8 gets closer to the pink one with a shift of about 0.25 µas, putting the deviations on $a_{1}$ and $\epsilon$ on a more similar footing. For $r_{\mathrm{inner}}=3M$, the upper part of the Schwarzschild green band is encompassed. Hence, the position of the inner edge of the accretion disc could be an important source of uncertainty, especially for the interpretation of the direct image. Nevertheless, for MAD discs, the presence of dynamically important near-horizon magnetic fields ensures that the plasma moves slowly all the way down to the BH horizon (Narayan et al. 2003; The Event Horizon Telescope Collaboration 2021), thus taking $r_{\mathrm{inner}}=r_{\mathrm{H}}$ is a fair choice.
C.5 Inclination
We analyse here the contribution of a strong inclination, namely $i=100^{\circ}$, to the study of the degeneracy on the peaks of the 1D intensity profiles. Although this viewing angle is not applicable to M87* or Sgr A* (The Event Horizon Telescope Collaboration 2022), the investigations of this paragraph could be useful when, with the future interferometric BHEX observations, a broader number of supermassive black holes will be within reach (Johnson et al. 2024). To properly account for the deformation due to projection effects, extensively discussed in Appendix A, the edges of the first lensing band are ray-traced on 100 polar angles on the screen instead of 14 as in Section 5; on the other hand, the radial resolution is unchanged. It is interesting to notice that, contrary to the low-inclination framework treated in the main text, here there are some privileged polar directions on the screen to look for metric deviations. In Figure C.4, the three bands are clearly separated on the left side, while they are widely superposed on the right. As it is quite evident from Figure C.4, the fractional asymmetry is higher than for $i=163^{\circ}$, reaching almost 40% for the primary image and 3% for the first photon ring. Those values are nevertheless degenerate between the various spacetimes as well as those of the size ratios.
Figure 16: Positions of the $n=1$ peaks, observed at $i=100^{\circ}$, in the case of a Schwarzschild BH (green), a RZ BH with $a_{1}|_{\,\epsilon=0}=0.1$ (pink) and a RZ BH with $\epsilon\,|_{\,a_{1}=0}=0.1$ (violet). Critical curves are dashed lines.
C.6 $a_{0}$ and $b_{0}$
The zeroth-order parameters of the RZ metric of Section 3 are directly related to the PPN parameters $\beta$ and $\gamma$ via the following equations, where $\epsilon\coloneqq 2M/r_{\mathrm{H}}-1$ (Rezzolla & Zhidenko 2014):
$a_{0}=\frac{(\beta-\gamma)(1+\epsilon)^{2}}{2}\quad\text{and}\quad b_{0}=\frac{(\gamma-1)(1+\epsilon)}{2}\>.$ (C5)
The observational constraints on the PPN parameters from solar system measurements (Will 2014)
$|\beta-1|\leq 2.3\times 10^{-4}\quad\text{and}\quad|\gamma-1|\leq 2.3\times 10^{-5}\>,$ (C6)
force then $a_{0}$ and $b_{0}$ to be small: $|a_{0}|\leq 10^{-4}$ and $|b_{0}|\leq 10^{-4}$. Moreover, the values of $\beta$ and $\gamma$, and therefore those of $a_{0}$ and $b_{0}$, are corroborated by stellar astrometry results in the galactic centre (GRAVITY collaboration 2020, 2024). Since the galactic centre also hosts a supermassive black hole, assuming that the PPN constraints hold in the surroundings of M87* too, as we did in the paper, is a reasonable choice.
Figure 17: Positions of the $n=1$ peaks for all the possible emission configurations in the case of a Schwarzschild black hole (green), a RZ black hole with non-vanishing parameter $b_{1}=0.1$ (blue) and a RZ black hole with non-vanishing parameters $a_{0},\,b_{0}=10^{-4}$ (yellow) or $a_{0},\,b_{0}=0.1$ (red). Critical curves are represented as dashed lines: the green and blue ones are superposed while the yellow one is slightly above them.
If PPN constraints are satisfied, fixing $a_{0}$ and $b_{0}$ to zero, as we have chosen to do in this article, is legit and useful to analyse the impact of more influential parameters. In fact, as can be seen from the orange bands of Figure C.5, comprising the positions of the intensity peaks for all the emission profiles of Table 2, a spacetime defined by the maximum allowed values of $a_{0}$ and $b_{0}$ from stellar system observations, that is, $a_{0}=b_{0}=10^{-4}$ with all other parameters vanishing, produces indiscernible differences from the Schwarzschild green case. If PPN constraints are not satisfied around M87*, which is possible in theories of gravity without a Birkhoff-like uniqueness theorem, the $a_{0}$ parameter is a relevant deviation to consider as it leads to observables comparable to those produced by $a_{1}$, see Figure C.5.
C.7 $b_{1}$
We repeat here the same study on the peaks positions for small deviations in the parameter $b_{1}$. Particularly, we choose $b_{1}=0.1$, similarly to what have been done for $\epsilon$ or $a_{1}$, and we fix all the other metric parameters to their Schwarzschild value. The results of this spacetime configuration are contained in the blue bands of Figure C.5, which does not significantly stand out from the green Schwarzschild band. In other words, small geometrical deviations in $b_{1}$ are degenerate with the astrophysical uncertainty both for the primary image and for the first-lensed photon ring, because $b_{i}$ parameters, with $i\in\mathbb{N}$, do not affect any of the spacetime properties related to the key image features, as explained in Section 6. Hence, this justifies our decision to ignore its effects in the main exploration of this article. We note that even for the extreme value $b_{1}=1$, the degeneracy is never broken.
Figure 18: Widths of the $n=0$ (left panels) and $n=1$ (right panels) images for all the possible emission configurations in the case of a Schwarzschild black hole (green), a RZ black hole with $a_{1}|_{\,\epsilon=0}=0.1$ (pink) and a RZ black hole with $\epsilon\,|_{\,a_{1}=0}=0.1$ (violet). The velocity of the disc is taken to be Keplerian for the upper panels, radial for the central panels and mixed for the lower panels.
Figure 19: Positions of peaks and widths of the $n=0$ (odd panels) and $n=1$ (even panels) profiles for all the possible emission configurations in the case of a Schwarzschild black hole (green), a RZ black hole with $a_{1}|_{\,\epsilon=0}=1.0$ (pink) and a RZ black hole with $\epsilon\,|_{\,a_{1}=0}=0.5$ (violet). The velocity of the disc is taken to be Keplerian for first row, radial for second row and mixed for the third row. Critical curves are represented as dashed lines.
C.8 $a_{1}$ and $\epsilon$
For small geometric deviations in $\epsilon$ and $a_{1}$, the width of the first photon ring is not an observable permitting to detect deviations from the Schwarzschild geometry, as can be seen in Figure C.6. However, for higher values of these same parameters, the width can become a complementary probe to the peak positions, although the latter remains a more robust probe. These statements can be visualised in Figures C.6 and C.7 where we represent the two extreme positive values allowed by conditions (3.2), namely $a_{1}|_{\,\epsilon=0}=1$ and $\epsilon\,|_{\,a_{1}=0}=0.5$. As already discussed in the main text, larger deviations require less stringent $M/D$ precisions to be observed. We remark also that these extreme deviations lead to more striking differences in the extent, but not the shape, of the bands which is a feature linked to the size of the corresponding lensing bands, which, despite not being observable, must contain the photon rings entirely (Cárdenas-Avendaño & Held 2024). Even for extreme RZ parametric deviations, the astrophysical-driven degeneracy on the image asymmetry and images size ratios introduced in Section 8.2 is never broken. Thus, relative measurements can never be used to distinguish a Schwarzschild black hole from a Rezzolla-Zhidenko spacetime.